### v2

# Low-Rank Tensor Decomposition over Finite Fields

Abstract
We show that finding rank-1, rank-2, and rank-3 decompositions of a 3D tensor over a fixed finite field can be done in polynomial time. However, if some cells in the tensor are allowed to have arbitrary values, then rank-2 is NP-hard over the integers modulo 2. We also explore rank-1 decomposition of a 3D tensor and of a matrix where some cells are allowed to have arbitrary values.
1 Introduction
Given a $p\times q\times s$ tensor (multidimensional array) $T$, a rank-$R$ decomposition is a collection of variables $A_{a,b},B_{c,d},C_{e,f}$ such that
$T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k.$
This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the slice $[M_{r,j}]_{j}$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
Tensor decomposition is at the heart of fast matrix multiplication, a problem that is the primary bottleneck of numerous linear algebra and graph combinatorics algorithms, such as matrix inversion and triangle detection. All asymptotically fast algorithms for matrix multiplication use a divide-and-conquer scheme, and finding an efficient scheme is equivalent to finding a decomposition of a certain tensor with low rank [2].
While significant progress has been made on this problem over the decades, most of the asymptotically fastest algorithms discovered use indirect constructions and convergence arguments that result in enormously large divide-and-conquer schemes, making the constant factors for these algorithms far too big for them to be useful in practice. For this reason, there has also been interest in finding explicit decompositions for the tensors of small divide-and-conquer schemes. With the introduction of AlphaTensor [3] and flip graphs [5] just over a year ago, several new low-rank decompositions have been discovered, especially over finite fields of characteristic 2 [1].
Decompositions using only integers are especially preferred, since they avoid having to use floating point arithmetic, and most decompositions of small divide-and-conquer schemes that have been found have happened to use only integers. This makes it useful to search for decompositions over finite fields (e.g., integers modulo 2), as doing so makes the search space finite and allows one to rule out large classes of integer solutions.
It is known that deciding whether a given tensor has a decomposition of rank $\leq R$ is NP-hard over finite fields via a reduction from 3SAT [4]. However, since we only want low-rank decompositions, we are more interested in hardness results for small fixed $R$.
1.1 Results
Our main result in this report is a polynomial-time algorithm for 3D tensor decomposition with rank $\leq 3$. For simplicity of asymptotic analysis, we focus on $n\times n\times n$ tensors $T$, although our algorithm easily generalizes to non-cube tensors.
Theorem 1.
For $R\leq 3$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for $f$ polynomial in $|\mathbb{F}|$ and exponential in $R$.
However, when allowing certain cells of the target tensor to have arbitrary value (“wildcards”), we have a hardness result over the integers mod 2 via a reduction from Not-All-Equal 3SAT (NAE-3SAT):
Theorem 2.
If some cells of a $p\times q\times s$ tensor $T$ are allowed to have arbitrary value, then determining if $T$ has a rank-2 decomposition over $\mathbb{Z}/2\mathbb{Z}$ is NP-hard.
We also sketch out polynomial-time algorithms for rank-1 decomposition with wildcards over a 3D tensor and a matrix, over an arbitrary fixed finite field.
2 Without wildcards
In this section we describe our algorithm for Theorem 1. The following observations of rank-$\leq 2$ decomposition over $\mathbb{Z}/2\mathbb{Z}$ serve as inspiration:
  • Rank-1 is easy since $abc=1$ forces $a=1$, $b=1$, and $c=1$, so $T=a\times b\times c$ if and only if $(a_{i}=\lor_{j,k}T_{i,j,k},\ b_{j}=\lor_{i,k}T_{i,j,k},\ c_{k}=\lor_{i,j}T_{i,j,k})$ is a solution.
  • Rank-2 at first glance is not so clear, but we notice that if $T=a_{0}\times b_{0}\times c_{0}+a_{1}\times b_{1}\times c_{1}$, then all slices $T_{i,:,:}$ are of the form $u(b_{0}\times c_{0})+v(b_{1}\times c_{1})$ for $u,v\in\mathbb{Z}/2\mathbb{Z}$, so each $T_{i,:,:}$ can only be one of four possibilities. These slices also have rank at most 2, so one could maybe do exhaustive casework over the ranks of these slices to find a rank-2 decomposition of $T$ (or prove that one does not exist).
Our algorithm generalizes to rank-3 and arbitrary finite fields. Suppose we want to find a rank-$R$ decomposition of the $n\times n\times n$ tensor $T$ over the finite field $\mathbb{F}$, i.e., $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k$. Our main insight is to fix every possible $A$ and then solve for $B,C$ in a clever manner. To ensure we do not enumerate too many $A$ (specifically, the number of possible $A$ we search over is $O(1)$ w.r.t. $n$), we extract a basis on the slices $T_{i,:,:}$.
Flatten each $T_{i,:,:}$ to form the $n\times n^{2}$ matrix $\overline{T}:=\left[T_{i,:,:}\texttt{.flatten}()\right]_{i}$111 We use numpy.ndarray.flatten to flatten an $n\times n$ matrix into a $n^{2}$-length vector, with elements in row-major order. For example, $\begin{bmatrix}a&b\\ c&d\end{bmatrix}\texttt{.flatten}()=\begin{bmatrix}a&b&c&d\end{bmatrix}$. ; run Gaussian elimination to get $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$, where rref denotes reduced row-echelon form and $M$ is invertible; let $J$ be the number of nonzero rows of $\mathrm{rref}\left(\overline{T}\right)$, which is equal to $\mathrm{rank}\left(\overline{T}\right)$. Then we can collect basis slices to form the tensor $T^{\prime}=\left[\mathrm{rref}\left(\overline{T}\right)_{j,:}\texttt{.reshape}(n,n)\right]_{0\leq j<J}$222 We use numpy.ndarray.reshape to reshape a $n^{2}$-length vector into a $n\times n$ matrix, with elements in row-major order. For example, $\begin{bmatrix}a&b&c&d\end{bmatrix}\texttt{.reshape}(2,2)=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$. , yielding $T_{i,:,:}=\sum_{0\leq j<J}M_{i,j}T^{\prime}_{j,:,:}$ and $T^{\prime}_{j,:,:}=\sum_{0\leq i<n}(M^{-1})_{j,i}T_{i,:,:}$. These relations allow us to convert between decompositions of $T$ and $T^{\prime}$:
$T^{\prime}=\sum_{r=0}^{R-1}U_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T=\sum_{r=0}^{R-1}\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{i}\times B_{r,:}\times C_{r,:}$
$T=\sum_{r=0}^{R-1}A_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T^{\prime}=\sum_{r=0}^{R-1}\left[\sum_{0\leq i<n}(M^{-1})_{j,i}A_{r,i}\right]_{i}\times B_{r,:}\times C_{r,:}$
Thus, finding a rank-$R$ decomposition of $T$ is equivalent to finding a rank-$R$ decomposition of $T^{\prime}$. We can then rewrite $T^{\prime}=\sum_{r}U_{r}\times B_{r}\times C_{r}$ into the block matrix form $\begin{bmatrix}T^{\prime}_{j,:,:}\end{bmatrix}_{0\leq j<J}=U^{\intercal}\begin{bmatrix}B_{r,:}\times C_{r,:}\end{bmatrix}_{r}$.
The upshot of constructing $T^{\prime}$ is that if there exists $A,B,C$ s.t. $T=\sum_{r}A_{r}\times B_{r,:}\times C_{r,:}$, then $\forall i,\ T_{i,:,:}\in\textrm{span}\left\{B_{r,:}\times C_{r,:}\right\}_{0\leq r<R}$, so $\left\{T_{i,:,:}\right\}_{0\leq i<n}$ spans a linear subspace of $\mathbb{F}^{n\times n}$ with rank at most $R$. This forces $T^{\prime}$ to have at most $R$ many slices, i.e., $J\leq R$, so there are at most $|\mathbb{F}|^{R^{2}}$ possible $U$ to fix. Thus, we avoid exponential dependence on $n$.
If $J=0$, then $T$ must be all-zero, so it trivially has a rank-$R$ decomposition (e.g., with all terms set to 0). Otherwise, for every possible $U$, run Gaussian elimination on $\Xi:=\left[\begin{array}[]{c|c}U^{\intercal}&\begin{bmatrix}\vdots\\ T^{\prime}_{j,:,:}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$ to obtain
$\Xi^{\prime}:=\left[\begin{array}[]{c|c}\Omega&\begin{bmatrix}\vdots\\ D_{j}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$, for $\Omega:=\mathrm{rref}\left(U^{\intercal}\right)$ and some $n\times n$ matrices $D_{j}$.
If there is some row $r$ s.t. $\Omega_{r,:}$ is all-zeros but $D_{r}$ is not, then there is no solution. Otherwise, since all all-zero rows of a rref matrix must be at the bottom, we can remove everything except the first $N$ rows of $\Xi^{\prime}$, where $N$ is the number of nonzero rows of $\Omega$.
Now we analyze all possible forms of the truncated rref matrix $E:=\Omega_{:N,:}$ with $R$ columns and $N\leq J\leq R$ rows, for $R\leq 3$:
  • For $R=1$, we must have $E=\begin{bmatrix}x\end{bmatrix}$, so we simply check whether $\mathrm{rank}\left(D_{0}\right)\leq\begin{cases}1&x\neq 0\\ 0&\textrm{else}\end{cases}$.
  • For $R=2$, we either have $E=\begin{bmatrix}1\\ &1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq 1\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 1$; or $E=\begin{bmatrix}1&x\end{bmatrix},\begin{bmatrix}0&1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$.
  • For $R=3$, there are a few more cases, of which one is nontrivial:
    • $E=\begin{bmatrix}1&x&y\end{bmatrix},\begin{bmatrix}0&1&x\end{bmatrix},\begin{bmatrix}0&0&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$;
    • $E=\begin{bmatrix}1\\ &1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq 1\ \forall i=0,1,2$;
    • $E=\begin{bmatrix}1&x\\ &&1\end{bmatrix},\begin{bmatrix}0&1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq(\#\textrm{ of nonzeros in }E_{i,:})\ \forall i=0,1$;
    • $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ (the nontrivial case):
      If $x=0$ or $y=0$, we can do the same thing as in $\begin{bmatrix}1&x\\ &&1\end{bmatrix}$ or $\begin{bmatrix}0&1\\ &&1\end{bmatrix}$, as the sets of nonzero terms of different rows in $E$ are pairwise disjoint.
      Otherwise, check whether $\exists i\textrm{ s.t. }\mathrm{rank}\left(D_{i}\right)=2$ (notice the $=$ instead of $\leq$ this time):
        if such $i$ exists, enumerate all rank-2 decompositions of $D_{i}=u_{0}\times v_{0}+u_{1}\times v_{1}$, and for each decomposition, check whether $\mathrm{rank}\left(D_{j}-u_{1}\times v_{1}\right)\leq 1$ for the other index $j\neq i$;
        otherwise, $\mathrm{rank}\left(D_{0}\right)\leq 1$ and $\mathrm{rank}\left(D_{1}\right)\leq 1$, so we can force $B_{2,:},C_{2,:}=\overrightarrow{0}$ and solve for $D_{0}=B_{0,:}\times C_{0,:},\ D_{1}=B_{1,:}\times C_{1,:}$;
In contrast to general tensors, rank decomposition of a $n\times n$ matrix (with the decomposition having smallest possible rank) can be computed in $O(n^{3})$ time via Gaussian elimination, so if all relevant checks in this casework pass, it is easy to construct a solution for $B,C$.
It remains to count how many rank-2 decompositions there are for a rank-2 matrix. Fortunately we have such a bound for arbitrary matrix rank which does not have exponential dependence on $n$:
Theorem 3.
Consider an $n\times m$ matrix $M$ over an arbitrary field with rank $r:=\mathrm{rank}\left(M\right)>0$. Let $F_{0}=\mathrm{rref}\left(M\right)_{:r,:}$ and $C_{0}$ satisfy $C_{0}F_{0}=M$. Then for any $n\times r$ matrix $C$ and $r\times m$ matrix $F$ s.t. $CF=M$, there is some invertible $r\times r$ $X$ s.t. $C=C_{0}X^{-1}$ and $F=XF_{0}$.
Proof.
First we note that $C_{0}$ exists and can be constructed via Gaussian elimination by maintaining the composition of all row operations in a separate matrix (and then truncating that matrix to its first $r$ columns). Additionally, $C_{0}$ is unique, because $\mathrm{rank}\left(F_{0}\right)=r$ and $r\leq\min\left(n,m\right)$ force the rows of $F_{0}$ to be linearly independent.
We have $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(CF\right)\subseteq\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)\leq r$, so $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(F\right)$. Since row reduction and removing all-zero rows preserves row-span, $\mathrm{rowspan}\left(F_{0}\right)=\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$.
We recall a property of rref matrices. Let $W$ be an arbitrary rref matrix with no all-zero rows. By definition of rref, all terms directly above and below any leading 1 must be 0, so in any linear combination of rows of $W$, leading 1s cannot be canceled out with any other terms. Thus, denoting $l_{i}$ as the column position of the leading 1 at row $i$, the position of the leftmost nonzero element of the row vector $\overrightarrow{v}:=\sum_{i}c_{i}W_{i,:}$ for scalars $c_{i}$ is $l_{k}$ for $k=\min{\left\{i:c_{i}\neq 0\right\}}$. Furthermore, $\overrightarrow{v}_{l_{i}}=c_{i}$, so given an arbitrary linear combination $\overrightarrow{v}$ of the rows of $W$ it is possible to determine the sequence of coefficients $c_{i}$ used to obtain the linear combination. Since the rows of $W$ are linearly independent, there is exactly one such sequence of coefficients $c_{i}$.
Since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, it is possible for a row vector in $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ to have its leftmost nonzero element be at any leading 1’s column in $F_{0}$, forcing $\mathrm{rref}\left(F\right)$ to have its set of column indices of leading 1s be the same as that of $F_{0}$, and thus forcing the leading 1s of $\mathrm{rref}\left(F\right)$ to be in the same set of locations as $\mathrm{rref}\left(F_{0}\right)$. Then, since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, and each row of $F_{0}$ has a 1 at the column index its own leading 1 and a 0 at all column indices of all other leading 1s, $\mathrm{rref}\left(F\right)$ itself must contain each row of $F_{0}$. Because leading 1s in a rref matrix must appear left-to-right as one goes from the top to the bottom row, this is enough to force $\mathrm{rref}\left(F\right)=F_{0}$.
By Gaussian elimination, there is some invertible $r\times r$ matrix $X$ s.t. $F=X\mathrm{rref}\left(F\right)=XF_{0}$, so $M=CF=CXF_{0}$; since the equation $M=AF_{0}$ has exactly one solution for $A$, $CX$ is forced to equal $C_{0}$, so $C=C_{0}X^{-1}$. ∎
The upshot of this result is that over a finite field $\mathbb{F}$, there are exactly $|\mathrm{GL}_{r}\left(\mathbb{F}\right)|=O(|\mathbb{F}|^{r^{2}})$ many rank-$r$ decompositions of a rank-$r$ matrix, and this quantity is independent of the number of rows or columns the matrix has. One can generate all of these decompositions by first constructing $(C_{0},F_{0})$ via Gaussian elimination in $O(n^{3})$ time, then iterate over every $r\times r$ matrix $X$, check whether $X$ is invertible in $O(r^{3})$ time (e.g., by checking if $\det\left(X\right)\neq 0$), then compute $(C,F)=(C_{0}X^{-1},F_{0}X^{-1})$ in $O((n+m)r^{2})$ time. Total running time is then $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{5})$. If a list of invertible $r\times r$ matrices and their inverses are precomputed, total running time is $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{2})$ instead.
In contrast, the number of rank-$(r+1)$ decompositions of $M$ might already be exponential in the dimensions of $M$ [^3] Consider the case where every element in $M$ is 1..
2.1 Runtime
In our algorithm, we first factor $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$ and construct $T^{\prime}$ via Gaussian elimination. Since $J=\mathrm{rank}\left(\overline{T}\right)$ must be at most $R$ in order for $T$ to have a rank-$R$ decomposition, we can terminate early if we detect that this condition fails, yielding a running time of $O(Rn^{3})$444 Given a $n\times m$ matrix $A$, we can construct $C,F$ s.t. $A=CF$, $F=\mathrm{rref}\left(A\right)$, and $C$ is invertible, as follows: 1. initialize $(C,F)=(I_{n},A)$ and $r=0$: 2. for $j$ from 00 to $m-1$ inclusive: (a) find some $i\geq r$ s.t. $F_{i,j}\neq 0$ (b) if such $i$ does not exist: skip to the next iteration of the for-loop (c) else: i. swap rows $i$ and $r$ of $F$, and swap columns $i$ and $r$ of $C$ ii. for $k=r+1,\dots n-1$: A. let $s=F_{k,j}$ B. set $F_{k,:}$ to $F_{k,:}-sF_{r,:}$ C. set $C_{:,r}$ to $C_{:,r}+sC_{:,k}$ (here $C_{:,r}$ denotes column $r$ of $C$) iii. increment $r$ The rank of $A$ is exactly the number of times we reach (c). If we only need to return $(C,F)$ under the condition $\mathrm{rank}\left(A\right)\leq R$, for some rank limit $R$, and return a failure statement otherwise, we can terminate early if we ever detect that (c) has been reached more than $R$ times. Thus, this procedure performs $O(Rn(m+n))$ operations. In our main algorithm, $A=\overline{T}$ has shape $n\times n^{2}$, so this procedure runs in $O(Rn^{3})$ time. .
Then we enumerate every possible $U$, of which there are $\leq|\mathbb{F}|^{JR}$ of them. For each $U$, we perform Gaussian elimination on the matrix $\Xi$, which has dimensions $J\times(R+1)$ and has each element in its rightmost column being a $n\times n$ matrix, so we incur $O\left(R^{3}n^{2}\right)$ runtime.
We then ensure there is no $r$ s.t. $\Omega_{r,:}=\overrightarrow{0}$ and $D_{r}\neq O$, which takes $O(J(R+n^{2}))$ time. If this check passes, we proceed to the rref case analysis, during which we compute the ranks of one or more matrices $D_{i}$, each of which takes $O(n^{3})$ time via Gaussian elimination.
If $R=3$ we also have to enumerate all rank-2 decompositions of a rank-2 $n\times n$ matrix $D_{i}$ in the worst case, which takes $O(n^{3}+|\mathbb{F}|^{4}n)$ time (if we precompute all invertible $2\times 2$ matrices over $\mathbb{F}$).
If we pass the case analysis, we construct $B$ and $C$ in $O(n^{3})$ time, then $A:=\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{r,i}$ in $O(RJn)$ time, and terminate.
Since we are guaranteed that $J\leq R$ (as otherwise no solution exists), our final runtime is
$O\left(Rn^{3}+\underbrace{|\mathbb{F}|^{R^{2}}}_{\textrm{fix }U}\left(\underbrace{R^{3}n^{2}}_{\textrm{row-reduce }\Xi}+\underbrace{R(n^{2}+R)}_{\textrm{check rows of }\Xi}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)+n^{3}+R^{2}n\right)\\ =O\left(|\mathbb{F}|^{R^{2}}\left(R^{3}n^{2}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)\right)$.
Thus, for any fixed $\mathbb{F}$ and fixed $R$ our algorithm runs in $O(n^{3})$ time, with hidden constant $f(|\mathbb{F}|,R)$ polynomial in $|\mathbb{F}|$ and exponential in $R$; for a loose bound we can set $f(|\mathbb{F}|,R)=|\mathbb{F}|^{R^{2}+4}R^{3}$. For $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ and $R=3$, the coefficient on the $n^{3}$ term is $\approx 2^{3^{2}}\cdot(3+1)=4\cdot 2^{9}=2048$.
2.2 Beyond rank 3
We do not know how to make our algorithm run in $\mathrm{poly}(n)$ time for $R=4$. Consider the case $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$, $\mathrm{rank}\left(D_{0}\right)=2,\ \mathrm{rank}\left(D_{1}\right)=2$: unlike when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$, we cannot simply force $B_{3,:},C_{3,:}=\overrightarrow{0}$, as the system of equations might become too restrictive. One might also notice that $D_{0}+D_{1}=B_{0,:}\times C_{0,:}+B_{1,:}\times C_{1,:}$, so $D_{0}+D_{1}$ has rank $\leq 2$; however, if it happens to have rank 1, then there still might be exponentially many solutions for $(B_{0,:},C_{0,:},B_{1,:},C_{1,:})$, so it is too expensive to simply iterate over all such solutions and check for feasibility of $(B_{2,:},C_{2,:},B_{3,:},C_{3,:})$.
3 With wildcards
When certain elements in $T$ are wildcards (i.e., allowed to have arbitrary value), tensor decomposition becomes much more difficult. In this section we prove Theorem 2, i.e., rank-2 decomposition with wildcards over $\mathbb{Z}/2\mathbb{Z}$ is NP-hard, via a reduction from NAE-3SAT.
Let $\mathcal{I}:=\begin{bmatrix}\begin{bmatrix}1&0\\ 0&0\end{bmatrix}&\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\end{bmatrix}$ be the $2\times 2\times 2$ tensor with 1s along the main diagonal and 0s everywhere else. It can be verified by brute force that the only rank-2 decompositions of $\mathcal{I}$ are $\begin{bmatrix}1\\ 0\end{bmatrix}^{\times 3}+\begin{bmatrix}0\\ 1\end{bmatrix}^{\times 3}$ and $\begin{bmatrix}0\\ 1\end{bmatrix}^{\times 3}+\begin{bmatrix}1\\ 0\end{bmatrix}^{\times 3}$ (even when allowing all-zero vectors). We can parameterize both of these solutions as $\begin{bmatrix}a\\ \neg a\end{bmatrix}^{\times 3}+\begin{bmatrix}\neg a\\ a\end{bmatrix}^{\times 3}$, yielding a variable gadget. By concatenating $n$ copies of $\mathcal{I}$ diagonally [^5] For an example, the result of concatenating $\begin{bmatrix}\begin{bmatrix}a_{0}&a_{1}\\ a_{2}&a_{3}\end{bmatrix}&\begin{bmatrix}a_{4}&a_{5}\\ a_{6}&a_{7}\end{bmatrix}\end{bmatrix}$ and $\begin{bmatrix}\begin{bmatrix}b_{0}&b_{1}\\ b_{2}&b_{3}\end{bmatrix}&\begin{bmatrix}b_{4}&b_{5}\\ b_{6}&b_{7}\end{bmatrix}\end{bmatrix}$ diagonally is $\begin{bmatrix}\begin{bmatrix}a_{0}&a_{1}\\ a_{2}&a_{3}\\ &&&\\ &&&\end{bmatrix}&\begin{bmatrix}a_{4}&a_{5}\\ a_{6}&a_{7}\\ &&&\\ &&&\end{bmatrix}&\begin{bmatrix}&&&\\ &&&\\ &&&\\ &&b_{0}&b_{1}\\ &&b_{2}&b_{3}\end{bmatrix}&\begin{bmatrix}&&&\\ &&&\\ &&&\\ &&b_{4}&b_{5}\\ &&b_{6}&b_{7}\end{bmatrix}\end{bmatrix},$ where the unmarked cells can be filled with arbitrary things. to form the target tensor $T$ and setting all other elements to be wildcards, we force all rank-2 decompositions of $T$ to be of the form [$\ v_{i}\\ \neg v_{i}\\ \vdots\end{bmatrix}^{\times 3}+\begin{bmatrix}\vdots\\ \neg v_{i}\\ v_{i}\\ \vdots\end{bmatrix}^{\times 3}$, for arbitrary variables $v_{i}$.
Since $\mathrm{NAE}\left(a,b,c\right)=\neg(abc\lor(\neg a)(\neg b)(\neg c))$, and the AND-clauses $abc$ and ($(\neg a)(\neg b)(\neg c)$ cannot both be equal at the same time, $\mathrm{NAE}\left(a,b,c\right)=\neg(abc\oplus(\neg a)(\neg b)(\neg c))=(abc+(\neg a)(\neg b)(\neg c)=0)$, where $\oplus$ denotes XOR and is isomorphic to addition mod 2. Because $T$ already forces all elements of vectors in the second outer product to be negations of vectors in the first outer product, we can encode $\mathrm{NAE}\left(v_{i},v_{j},v_{k}\right)$ by setting $T_{2i,2j,2k}=0$; for negative literals, we add 1 to the corresponding tensor index, e.g., $\mathrm{NAE}\left(v_{i},v_{j},\neg v_{k}\right)$ can be encoded by setting $T_{2i,2j,2k+1}=0$.
Thus, we convert a NAE-3SAT instance with $n$ variables and $m$ clauses into a rank-2 decomposition instance on a $2n\times 2n\times 2n$ tensor with $8n+m$ fixed elements and all other elements being wildcards.
Over arbitrary finite fields, there is an issue of scale invariance, i.e., each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t. $\alpha\beta\gamma=1$, allowing any one tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor. This scaling freedom makes constructing gadgets for NP-hardness less obvious.
3.1 Rank 1 with wildcards
Here we outline a polynomial-time algorithm for rank-1 decomposition with wildcards over a fixed finite field $\mathbb{F}$.
First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ and set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t. $a_{i}b_{j}c_{k}=T_{i,j,k}\ \forall(i,j,k)\in S$.
When $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$, $abc=1$ implies $(a,b,c)=(1,1,1)$, so a solution exists if and only if setting
$a_{i}=\lor_{j,k:(i,j,k)\in S}T_{i,j,k},\ b_{j}=\lor_{i,k:(i,j,k)\in S}T_{i,j,k},\ c_{k}=\lor_{i,j:(i,j,k)\in S}T_{i,j,k}$
yields a solution; this is because this construction produces the minimal set of elements in $a\times b\times c$ that are forced to be 1s by the values $T_{i,j,k}$, and this set covers all $(i,j,k)\in S$ that have $T_{i,j,k}=1$. Thus for $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ there is an algorithm for rank-1 wildcard decomposition that runs in $O(n^{3})$ time with low constant factor.
For an arbitrary fixed finite field $\mathbb{F}$, we outline a polynomial-time algorithm.
Let $S^{\prime}:=\left\{(i,j,k)\in S:T_{i,j,k}\neq 0\right\}$ denote all points with nonzero values and $I:=\left\{i:\exists(i,j,k)\in S^{\prime}\right\}$, $J:=\left\{j:\exists(i,j,k)\in S^{\prime}\right\}$, $K:=\left\{k:\exists(i,j,k)\in S^{\prime}\right\}$ denote the sets of coordinate values of the nonzero points along each dimension. Since $T_{i,j,k}\neq 0\Rightarrow a_{i},b_{j},c_{k}\neq 0$, we have that $a_{i}\neq 0\ \forall i\in I$, $b_{j}\neq 0\ \forall j\in J$, and $c_{k}\neq 0\ \forall k\in K$, so we check whether $T_{i,j,k}\neq 0\ \forall(i,j,k)\in(I\times J\times K)\cap S$. If this check fails, there is no solution.
Otherwise, we proceed with the observation that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve pure products, i.e., expressions of the form $\left(\prod_{i}a_{i}^{p_{i}}\right)\left(\prod_{j}b_{j}^{q_{j}}\right)\left(\prod_{k}c_{k}^{r_{k}}\right)$, for integers $p_{i},q_{j},r_{k}$. If we allow arbitrary multiplication and division between such expressions, all resulting expressions will still be of this form, although some $p_{i},q_{j},r_{k}$ might be negative.
This observation suggests that we can solve this problem somewhat like a linear system of equations: construct a $|S^{\prime}|\times 3n$ integer matrix $E$ and a $|S^{\prime}|\times 1$ $\mathbb{F}$-vector $W$ s.t. each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ for $(i,j,k)\in S^{\prime}$ is encoded as $\prod_{m}v_{m}^{E_{l,m}}=W_{l}$ for some $l$, where the variables have been re-indexed as $v_{0,\dots n-1}=a_{0,\dots n-1},\ v_{n,\dots 2n-1}=b_{0,\dots n-1},\ v_{2n,\dots 3n-1}=c_{0,\dots n-1}$. Since we restrict ourselves to $S^{\prime}$, each $W_{l}$ is nonzero. During row reductions, elements in $W$ have to be treated differently: when adding between two rows, multiply their values in $W$; when scaling a row by $s$, raise its value in $W$ to the $s$-th power.
However, once we reduce $\left[\begin{array}[]{c|c}E&W\end{array}\right]$ to some normal form, elements in $E$ might not be ±$\pm 1$, so extracting a solution for the variables $v_{i}$ may involve computing $r$-th roots in $\mathbb{F}$. An explicit example of this phenomenon is the tensor
$T=\begin{bmatrix}\begin{bmatrix}1&?\\ ?&1\end{bmatrix},\begin{bmatrix}?&2\\ 1&?\end{bmatrix}\end{bmatrix},$
where ? represents wildcards (equivalently, $S=\left\{(0,0,0),(1,1,0),(0,1,1),(1,0,1)\right\}$, $T_{0,0,0}=1,\ T_{1,1,0}=1,\ T_{0,1,1}=2,\ T_{1,0,1}=1$). Solving this instance is equivalent to solving the equations
$\left\{a_{0}b_{0}c_{0}=1,\ a_{1}b_{1}c_{0}=1,\ a_{0}b_{1}c_{1}=2,\ a_{1}b_{0}c_{1}=1\right\};$
by combining these equations, we get $\frac{(a_{0}b_{1}c_{1})(a_{1}b_{0}c_{1})}{(a_{0}b_{0}c_{0})(a_{1}b_{1}c_{0})}=\frac{2\cdot 1}{1\cdot 1}\Rightarrow\left(\frac{c_{1}}{c_{0}}\right)^{2}=2$, so solving this equation requires us to compute a square root of 2.
To resolve this issue, we use the fact that the multiplicative group of any finite field is cyclic, so if $\mathbb{F}$ is fixed we can find a primitive element (i.e., multiplicative generator) $\mu$ via brute force and define $u_{i},w_{i}$ s.t. $v_{i}=\mu^{u_{i}},\ W_{i}=\mu^{w_{i}}$; then our problem reduces to solving the linear system of equations $Eu=w$ for the vector $u$ over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
From [6], solving the matrix-vector equation $xA=b$ over $\mathbb{Z}/N\mathbb{Z}$ for $n\times m$ $A$ can be done in $O(nm^{\omega-1})$ many operations, where $\omega\leq 2.371552$ [8] is the exponent for matrix multiplication, with all integers throughout the algorithm having magnitude $\leq N$. By transposing our system to $u^{\intercal}E^{\intercal}=w^{\intercal}$ and noting that $E$ has dimensions $|S^{\prime}|\times 3n$, we see that solving our system takes $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$. For fixed $\mathbb{F}$, since all integers used in this algorithm have magnitude $\leq|\mathbb{F}|-1$, each operation takes $O(1)$ time, so our algorithm runs in polynomial time. If we account for the number of wildcards in $T$ in the time complexity, then by letting $k$ be the number of wildcards, $|S^{\prime}|\leq|S|=n^{3}-k$, so we get $O(n(n^{3}-k)^{\omega-1})$.
The main bottleneck in this algorithm is finding a primitive element of $\mathbb{F}$, for which an efficient algorithm is not known [7]. On the other hand, for $\mathbb{F}=\mathbb{F}_{p^{k}}$, the total input size of a problem instance would be linear w.r.t. $\log(p)$ and $k$ [^6] Every element in $\mathbb{F}_{p^{k}}$ is a linear combination of $1,\alpha,\dots,\alpha^{k-1}$ with coefficients in $\mathbb{Z}/p\mathbb{Z}$, for some element $\alpha$, so every element can be specified with a $k$-element vector over $\mathbb{Z}/p\mathbb{Z}$.. Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is w.r.t. input size when $p,k$ are not fixed parameters.
3.2 Rank-1 with wildcards for matrices
One may wonder whether matrix rank factorization with wildcards is NP-hard. The issue is that matrix factorizations have an immense degree of freedom, as $CF=M$ implies $(CX^{-1})(XF)=M$, so constructing gadgets seems very difficult even for rank $R=2$.
Additionally, for rank $R=1$, we sketch an $O(n^{3})$ algorithm that does not contain $|\mathbb{F}|$ in its asymptotic running time, in contrast to the analogous problem for 3D tensors. Again we formulate the problem as follows: given a subset $S\subseteq\left\{0,\dots n-1\right\}^{2}$ and set of elements $T_{i,j}\in\mathbb{F}\ \forall(i,j)\in S$, find variables $a_{0,\dots n-1},b_{0,\dots n-1}$ s.t. $a_{i}b_{j}=T_{i,j}$.
Like before, define $S^{\prime}:=\left\{(i,j)\in S:T_{i,j}\neq 0\right\}$, $I:=\left\{i:(i,j)\in S^{\prime}\right\}$, $J:=\left\{j:(i,j)\in S^{\prime}\right\}$. Then $a_{i}\neq 0\ \forall i\in I$ and $b_{j}\neq 0\ \forall j\in J$, so we check whether $T_{i,j}\neq 0\ \forall(i,j)\in(I\times J)\cap S$; if the check fails then there is no solution.
Otherwise, we notice that since every equation $a_{i}b_{j}=T_{i,j}$ constrains the product of two variables, knowing one variable immediately forces the value of the other variable. This motivates us to construct a graph $G$ with vertex set $S^{\prime}$ where there is an edge between every pair of vertices $(i,j),(i^{\prime},j^{\prime})$ s.t. $i=i^{\prime}$ or $j=j^{\prime}$.
Choose an arbitrary $(i_{0},j_{0})\in S^{\prime}$: WLOG we can set $a_{i_{0}}=1,b_{j_{0}}=T_{i_{0},j_{0}}$ because any solution $(a,b)$ to the original problem can be scaled to $(\alpha a,\frac{1}{\alpha}b)$ for any $\alpha\neq 0$. Then do a breadth-first search over $S^{\prime}$ starting from $(i_{0},j_{0})$: when we visit each new vertex $(i,j)\in S$, we will have already determined $a_{i}$ and $b_{j}$, so for each neighbor of the form $(i^{\prime},j)$ we set $a_{i^{\prime}}=T_{i^{\prime},j}/b_{j}$, and for each neighbor of the form $(i,j^{\prime})$ we set $b_{j^{\prime}}=T_{i,j^{\prime}}/a_{i}$. Each neighbor that was not already visited is added to a queue of vertices to visit. Note that because every $T_{i,j}$ we deal with is nonzero, we will never encounter division by 0 in this procedure.
If at any point we update a variable $a_{i}$ or $b_{j}$ that already has a assigned value and set it to a different value, then we have encountered an inconsistency and there is no solution to the original problem. Otherwise, at the end of this procedure we will have searched through the entire connected component of $G$ that contains $(i_{0},j_{0})$, so we do another breadth-first search starting at an arbitrary unvisited vertex in $G$, and repeat until all vertices in $G$ have been visited. If no inconsistencies were ever encountered, we can set all variables $a_{i}$ and $b_{j}$ that were not already assigned a value to 0, and we have a solution.
Since breadth-first search over a graph with $V$ vertices and $E$ edges takes $O(V+E)$ time, and $G$ has $O(n^{2})$ vertices and $O(n^{3})$ edges, our algorithm runs in $O(n^{3})$ time.
The key for correctness is that if two vertices $(i,j),(i^{\prime},j^{\prime})$ are in different connect components, then $(i^{\prime},j^{\prime})$ is not adjacent to any vertex $(x,y)$ in the connected component containing $(i,j)$, so $i^{\prime}\neq x$ and $j^{\prime}\neq y$. This means that $a_{i^{\prime}}$ and $b_{j^{\prime}}$ are not directly or indirectly involved in any equations with elements $T_{x,y}$ for each $(x,y)$ in the connected component of $(i,j)$. Thus, after we finish searching over each connected component, all remaining unassigned variables $a_{i}$ and $b_{j}$ are completely independent of all previously assigned variables, so as we search through the next connected component we will never encounter inconsistencies with variables assigned during a previous connected component. This independence between connected components also allows us to assume WLOG that $a_{i_{0}}=1$ and $b_{j_{0}}=T_{i_{0},j_{0}}$ at each individual connected component.
If we tried generalizing this algorithm to 3D tensors, we would encounter a problem: two vertices would have to match 2 out of 3 coordinates in order to be adjacent to each other, because each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ constrains the product of three variables instead of two, so two variables have to be known beforehand to force the value of the third one. But then if $(i,j,k)$ and $(i^{\prime},j^{\prime},k^{\prime})$ occupy different components, it is still possible for $(i^{\prime},j^{\prime},k^{\prime})$ to share 1 out of 3 coordinates with a vertex contained in the connected component containing $(i,j,k)$, so different connected components can still share common variables and thus interact with each other in a nontrivial manner.
4 Future directions
We conclude with some open questions:
  • For what $R\geq 4$ can rank-$R$ tensor decomposition (without wildcards) be computed in polynomial time over a fixed finite field?
  • For $R\leq 3$, how low can the constant factor of rank-$R$ decomposition be w.r.t. to the size of the finite field and w.r.t. $R$?
  • What is the parameterized complexity of rank-2 wildcard decomposition w.r.t. the number of wildcard elements in the target tensor? On a related note, how asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
  • What is the parameterized complexity of rank-1 wildcard decomposition w.r.t. the size or the characteristic of the underlying finite field?
  • Which of our algorithms be generalized beyond finite fields, such as the integers?
Acknowledgments
We thank Prof. Erik Demaine and TAs Josh Brunner, Lily Chung, and Jenny Diomidova for running MIT 6.5440 Fall 2023, which made this project possible. We also thank them for giving feedback on an earlier version of this report, some of which were requests for certain details and open questions to be addressed, including rank-1 wildcard decomposition.

### v3

@@
 Abstract
-We show that finding rank-1, rank-2, and rank-3 decompositions of a 3D tensor over a fixed finite field can be done in polynomial time.
+We show that finding rank-$R$ decompositions of a 3D tensor, for $R\leq 4$, over a fixed finite field can be done in polynomial time.
 However, if some cells in the tensor are allowed to have arbitrary values, then rank-2 is NP-hard over the integers modulo 2.
@@
 Theorem 1.
-For $R\leq 3$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for $f$ polynomial in $|\mathbb{F}|$ and exponential in $R$.
+For $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for some function $f$.
+We show tighter runtime analysis in Section 2.3.1.
 However, when allowing certain cells of the target tensor to have arbitrary value (“wildcards”), we have a hardness result over the integers mod 2 via a reduction from Not-All-Equal 3SAT (NAE-3SAT):
@@
 We also sketch out polynomial-time algorithms for rank-1 decomposition with wildcards over a 3D tensor and a matrix, over an arbitrary fixed finite field.
+1.2 Notation
+  • We use NumPy notation for indexing on tensors.
+    • e.g. for a matrix $M$:
+        $M_{i,:}$ denotes the $i$-th row
+        $M_{:i,:}$ denotes the matrix truncated to its first $i$ rows
+        $M_{:,i}$ denotes the $i$-th column
+  • All tensors and sequences are 0-indexed.
+  • The reduced row echelon form of a matrix $M$ is denoted as $\mathrm{rref}\left(M\right)$.
 2 Without wildcards
-In this section we describe our algorithm for Theorem 1. The following observations of rank-$\leq 2$ decomposition over $\mathbb{Z}/2\mathbb{Z}$ serve as inspiration:
-  • Rank-1 is easy since $abc=1$ forces $a=1$, $b=1$, and $c=1$, so $T=a\times b\times c$ if and only if $(a_{i}=\lor_{j,k}T_{i,j,k},\ b_{j}=\lor_{i,k}T_{i,j,k},\ c_{k}=\lor_{i,j}T_{i,j,k})$ is a solution.
-  • Rank-2 at first glance is not so clear, but we notice that if $T=a_{0}\times b_{0}\times c_{0}+a_{1}\times b_{1}\times c_{1}$, then all slices $T_{i,:,:}$ are of the form $u(b_{0}\times c_{0})+v(b_{1}\times c_{1})$ for $u,v\in\mathbb{Z}/2\mathbb{Z}$, so each $T_{i,:,:}$ can only be one of four possibilities.
-These slices also have rank at most 2, so one could maybe do exhaustive casework over the ranks of these slices to find a rank-2 decomposition of $T$ (or prove that one does not exist).
-Our algorithm generalizes to rank-3 and arbitrary finite fields.
-Suppose we want to find a rank-$R$ decomposition of the $n\times n\times n$ tensor $T$ over the finite field $\mathbb{F}$, i.e., $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k$.
-Our main insight is to fix every possible $A$ and then solve for $B,C$ in a clever manner.
-To ensure we do not enumerate too many $A$ (specifically, the number of possible $A$ we search over is $O(1)$ w.r.t.
-$n$), we extract a basis on the slices $T_{i,:,:}$.
-Flatten each $T_{i,:,:}$ to form the $n\times n^{2}$ matrix $\overline{T}:=\left[T_{i,:,:}\texttt{.flatten}()\right]_{i}$111 We use numpy.ndarray.flatten to flatten an $n\times n$ matrix into a $n^{2}$-length vector, with elements in row-major order.
-For example, $\begin{bmatrix}a&b\\ c&d\end{bmatrix}\texttt{.flatten}()=\begin{bmatrix}a&b&c&d\end{bmatrix}$.
-; run Gaussian elimination to get $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$, where rref denotes reduced row-echelon form and $M$ is invertible; let $J$ be the number of nonzero rows of $\mathrm{rref}\left(\overline{T}\right)$, which is equal to $\mathrm{rank}\left(\overline{T}\right)$.
-Then we can collect basis slices to form the tensor $T^{\prime}=\left[\mathrm{rref}\left(\overline{T}\right)_{j,:}\texttt{.reshape}(n,n)\right]_{0\leq j<J}$222 We use numpy.ndarray.reshape to reshape a $n^{2}$-length vector into a $n\times n$ matrix, with elements in row-major order.
-For example, $\begin{bmatrix}a&b&c&d\end{bmatrix}\texttt{.reshape}(2,2)=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$.
-, yielding $T_{i,:,:}=\sum_{0\leq j<J}M_{i,j}T^{\prime}_{j,:,:}$ and $T^{\prime}_{j,:,:}=\sum_{0\leq i<n}(M^{-1})_{j,i}T_{i,:,:}$.
-These relations allow us to convert between decompositions of $T$ and $T^{\prime}$:
-$T^{\prime}=\sum_{r=0}^{R-1}U_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T=\sum_{r=0}^{R-1}\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{i}\times B_{r,:}\times C_{r,:}$
-$T=\sum_{r=0}^{R-1}A_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T^{\prime}=\sum_{r=0}^{R-1}\left[\sum_{0\leq i<n}(M^{-1})_{j,i}A_{r,i}\right]_{i}\times B_{r,:}\times C_{r,:}$
-Thus, finding a rank-$R$ decomposition of $T$ is equivalent to finding a rank-$R$ decomposition of $T^{\prime}$.
-We can then rewrite $T^{\prime}=\sum_{r}U_{r}\times B_{r}\times C_{r}$ into the block matrix form $\begin{bmatrix}T^{\prime}_{j,:,:}\end{bmatrix}_{0\leq j<J}=U^{\intercal}\begin{bmatrix}B_{r,:}\times C_{r,:}\end{bmatrix}_{r}$.
-The upshot of constructing $T^{\prime}$ is that if there exists $A,B,C$ s.t.
-$T=\sum_{r}A_{r}\times B_{r,:}\times C_{r,:}$, then $\forall i,\ T_{i,:,:}\in\textrm{span}\left\{B_{r,:}\times C_{r,:}\right\}_{0\leq r<R}$, so $\left\{T_{i,:,:}\right\}_{0\leq i<n}$ spans a linear subspace of $\mathbb{F}^{n\times n}$ with rank at most $R$.
-This forces $T^{\prime}$ to have at most $R$ many slices, i.e., $J\leq R$, so there are at most $|\mathbb{F}|^{R^{2}}$ possible $U$ to fix.
-Thus, we avoid exponential dependence on $n$.
-If $J=0$, then $T$ must be all-zero, so it trivially has a rank-$R$ decomposition (e.g., with all terms set to 0).
-Otherwise, for every possible $U$, run Gaussian elimination on $\Xi:=\left[\begin{array}[]{c|c}U^{\intercal}&\begin{bmatrix}\vdots\\ T^{\prime}_{j,:,:}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$ to obtain
-$\Xi^{\prime}:=\left[\begin{array}[]{c|c}\Omega&\begin{bmatrix}\vdots\\ D_{j}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$, for $\Omega:=\mathrm{rref}\left(U^{\intercal}\right)$ and some $n\times n$ matrices $D_{j}$.
-If there is some row $r$ s.t.
-$\Omega_{r,:}$ is all-zeros but $D_{r}$ is not, then there is no solution.
-Otherwise, since all all-zero rows of a rref matrix must be at the bottom, we can remove everything except the first $N$ rows of $\Xi^{\prime}$, where $N$ is the number of nonzero rows of $\Omega$.
-Now we analyze all possible forms of the truncated rref matrix $E:=\Omega_{:N,:}$ with $R$ columns and $N\leq J\leq R$ rows, for $R\leq 3$:
-  • For $R=1$, we must have $E=\begin{bmatrix}x\end{bmatrix}$, so we simply check whether $\mathrm{rank}\left(D_{0}\right)\leq\begin{cases}1&x\neq 0\\ 0&\textrm{else}\end{cases}$.
-  • For $R=2$, we either have $E=\begin{bmatrix}1\\ &1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq 1\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 1$; or $E=\begin{bmatrix}1&x\end{bmatrix},\begin{bmatrix}0&1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$.
-  • For $R=3$, there are a few more cases, of which one is nontrivial:
-    • $E=\begin{bmatrix}1&x&y\end{bmatrix},\begin{bmatrix}0&1&x\end{bmatrix},\begin{bmatrix}0&0&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$;
-    • $E=\begin{bmatrix}1\\ &1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq 1\ \forall i=0,1,2$;
-    • $E=\begin{bmatrix}1&x\\ &&1\end{bmatrix},\begin{bmatrix}0&1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq(\#\textrm{ of nonzeros in }E_{i,:})\ \forall i=0,1$;
-    • $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ (the nontrivial case):
-      If $x=0$ or $y=0$, we can do the same thing as in $\begin{bmatrix}1&x\\ &&1\end{bmatrix}$ or $\begin{bmatrix}0&1\\ &&1\end{bmatrix}$, as the sets of nonzero terms of different rows in $E$ are pairwise disjoint.
-      Otherwise, check whether $\exists i\textrm{ s.t. }\mathrm{rank}\left(D_{i}\right)=2$ (notice the $=$ instead of $\leq$ this time):
-        if such $i$ exists, enumerate all rank-2 decompositions of $D_{i}=u_{0}\times v_{0}+u_{1}\times v_{1}$, and for each decomposition, check whether $\mathrm{rank}\left(D_{j}-u_{1}\times v_{1}\right)\leq 1$ for the other index $j\neq i$;
-        otherwise, $\mathrm{rank}\left(D_{0}\right)\leq 1$ and $\mathrm{rank}\left(D_{1}\right)\leq 1$, so we can force $B_{2,:},C_{2,:}=\overrightarrow{0}$ and solve for $D_{0}=B_{0,:}\times C_{0,:},\ D_{1}=B_{1,:}\times C_{1,:}$;
-In contrast to general tensors, rank decomposition of a $n\times n$ matrix (with the decomposition having smallest possible rank) can be computed in $O(n^{3})$ time via Gaussian elimination, so if all relevant checks in this casework pass, it is easy to construct a solution for $B,C$.
-It remains to count how many rank-2 decompositions there are for a rank-2 matrix. Fortunately we have such a bound for arbitrary matrix rank which does not have exponential dependence on $n$:
-Theorem 3.
-Consider an $n\times m$ matrix $M$ over an arbitrary field with rank $r:=\mathrm{rank}\left(M\right)>0$.
-Let $F_{0}=\mathrm{rref}\left(M\right)_{:r,:}$ and $C_{0}$ satisfy $C_{0}F_{0}=M$.
-Then for any $n\times r$ matrix $C$ and $r\times m$ matrix $F$ s.t.
-$CF=M$, there is some invertible $r\times r$ $X$ s.t.
-$C=C_{0}X^{-1}$ and $F=XF_{0}$.
+2.1 Preliminary results about matrix factorizations
+Before describing our algorithm for Theorem 1, we present several lemmas on matrix rank decompositions that we will use extensively.
+The first lemma is a direct result of Gaussian elimination:
+Lemma 1.
+Given any matrix $M\in\mathbb{F}^{m\times n}$, $\exists T\in\mathrm{GL}\left(m,\ \mathbb{F}\right)$ s.t. $TM=\mathrm{rref}\left(M\right)$.
+It follows that inverting a $n\times n$ matrix can be done in $O(n^{3})$ time.
+The next lemma is the uniqueness of rref (up to all-zero rows):
+Lemma 2.
+Two rref matrices with the same rowspan must be equal, up to the presence of all-zero rows.
 Proof.
-First we note that $C_{0}$ exists and can be constructed via Gaussian elimination by maintaining the composition of all row operations in a separate matrix (and then truncating that matrix to its first $r$ columns).
-Additionally, $C_{0}$ is unique, because $\mathrm{rank}\left(F_{0}\right)=r$ and $r\leq\min\left(n,m\right)$ force the rows of $F_{0}$ to be linearly independent.
-We have $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(CF\right)\subseteq\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)\leq r$, so $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(F\right)$.
-Since row reduction and removing all-zero rows preserves row-span, $\mathrm{rowspan}\left(F_{0}\right)=\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$.
-We recall a property of rref matrices.
-Let $W$ be an arbitrary rref matrix with no all-zero rows.
-By definition of rref, all terms directly above and below any leading 1 must be 0, so in any linear combination of rows of $W$, leading 1s cannot be canceled out with any other terms.
-Thus, denoting $l_{i}$ as the column position of the leading 1 at row $i$, the position of the leftmost nonzero element of the row vector $\overrightarrow{v}:=\sum_{i}c_{i}W_{i,:}$ for scalars $c_{i}$ is $l_{k}$ for $k=\min{\left\{i:c_{i}\neq 0\right\}}$.
-Furthermore, $\overrightarrow{v}_{l_{i}}=c_{i}$, so given an arbitrary linear combination $\overrightarrow{v}$ of the rows of $W$ it is possible to determine the sequence of coefficients $c_{i}$ used to obtain the linear combination.
-Since the rows of $W$ are linearly independent, there is exactly one such sequence of coefficients $c_{i}$.
-Since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, it is possible for a row vector in $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ to have its leftmost nonzero element be at any leading 1’s column in $F_{0}$, forcing $\mathrm{rref}\left(F\right)$ to have its set of column indices of leading 1s be the same as that of $F_{0}$, and thus forcing the leading 1s of $\mathrm{rref}\left(F\right)$ to be in the same set of locations as $\mathrm{rref}\left(F_{0}\right)$.
-Then, since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, and each row of $F_{0}$ has a 1 at the column index its own leading 1 and a 0 at all column indices of all other leading 1s, $\mathrm{rref}\left(F\right)$ itself must contain each row of $F_{0}$.
-Because leading 1s in a rref matrix must appear left-to-right as one goes from the top to the bottom row, this is enough to force $\mathrm{rref}\left(F\right)=F_{0}$.
-By Gaussian elimination, there is some invertible $r\times r$ matrix $X$ s.t.
-$F=X\mathrm{rref}\left(F\right)=XF_{0}$, so $M=CF=CXF_{0}$; since the equation $M=AF_{0}$ has exactly one solution for $A$, $CX$ is forced to equal $C_{0}$, so $C=C_{0}X^{-1}$.
+Consider an rref matrix $W=\left[\begin{array}[]{c}w_{0}\\ \hline\cr\vdots\\ \hline\cr w_{k-1}\\ \hline\cr O\end{array}\right]$, where each $w_{i}\neq\overrightarrow{0}$ has its leading 1 at column $\ell_{i}$, with $\ell_{0}<\dots<\ell_{k-1}$.
+Consider a linear combination of the rows, $v:=\sum_{i}\alpha_{i}w_{i}$, where the $\alpha_{i}$ are not all zero: then $v_{\ell_{i}}=\alpha_{i}$, since by definition of rref, every column of $W$ that has a leading 1 has all of its other elements set to 0.
+As a consequence, if we are instead given some $v$ that is known to be a linear combination of $w_{i}$, then the list of coefficients $\alpha_{i}$ of that linear combination are unique and can immediately be deduced by reading each $v_{\ell_{i}}$ element.
+Additionally, the leading nonzero term of $v$ must be at some $\ell_{i}$-th column, as we can use the following casework: if $\alpha_{0}\neq 0$, the leading nonzero term of $v$ is $\ell_{0}$, since each $w_{i}$ is all zeros to the left of its leading 1: else if $\alpha_{1}\neq 0$, the leading nonzero term of $v$ is $\ell_{1}$: etc.
+Now suppose there is another rref matrix $W^{\prime}$ s.t.
+$\mathrm{rowspan}\left(W\right)=\mathrm{rowspan}\left(W^{\prime}\right)$: then $W^{\prime}$ cannot have a leading 1 at a column that is not one of the $\ell_{i}$-th, otherwise $\mathrm{rowspan}\left(W^{\prime}\right)$ would contain a row-vector that is not in $\mathrm{rowspan}\left(W\right)$.
+Conversely, $W^{\prime}$ must contain a leading 1 at every $\ell_{i}$-th column, since $w_{i}\in\mathrm{rowspan}\left(W\right)$ for each $i$.
+Thus, for each $0\leq i<k$, $W^{\prime}_{i,:}$ has a leading 1 at column $\ell_{i}$.
+Finally, since $w_{i}\in\mathrm{rowspan}\left(W^{\prime}\right)$ and, by definition of rref, $w_{i}$ contains a 1 at column $\ell_{i}$ and a 0 at all other leading columns $\ell_{j}$ for $j\neq i$, the only way for $w_{i}$ to be a linear combination of the rows of $W^{\prime}$ is if $w_{i}=W^{\prime}_{i,:}$.
+Thus, $W_{:k,:}=W^{\prime}_{:k,:}$.
 ∎
-The upshot of this result is that over a finite field $\mathbb{F}$, there are exactly $|\mathrm{GL}_{r}\left(\mathbb{F}\right)|=O(|\mathbb{F}|^{r^{2}})$ many rank-$r$ decompositions of a rank-$r$ matrix, and this quantity is independent of the number of rows or columns the matrix has.
-One can generate all of these decompositions by first constructing $(C_{0},F_{0})$ via Gaussian elimination in $O(n^{3})$ time, then iterate over every $r\times r$ matrix $X$, check whether $X$ is invertible in $O(r^{3})$ time (e.g., by checking if $\det\left(X\right)\neq 0$), then compute $(C,F)=(C_{0}X^{-1},F_{0}X^{-1})$ in $O((n+m)r^{2})$ time.
-Total running time is then $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{5})$.
-If a list of invertible $r\times r$ matrices and their inverses are precomputed, total running time is $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{2})$ instead.
-In contrast, the number of rank-$(r+1)$ decompositions of $M$ might already be exponential in the dimensions of $M$ [^3] Consider the case where every element in $M$ is 1..
+The previous two lemmas yield an important result about full-rank factorizations, i.e. rank-$r$ factorizations of a rank-$r$ matrix:
+Lemma 3.
+Consider an $m\times n$ matrix $M$ over an arbitrary field with rank $r:=\mathrm{rank}\left(M\right)>0$.
+Then there exists $C_{0}\in\mathbb{F}^{m\times r},\ F_{0}\in\mathbb{F}^{r\times n}$ s.t.
+$C_{0}F_{0}=M$.
+Furthermore, for any $C\in\mathbb{F}^{m\times r},\ F\in\mathbb{F}^{r\times n}$ s.t.
+$CF=M$, $\exists X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$ s.t.
+$C=C_{0}X$ and $F=X^{-1}F_{0}$.
+Proof.
+By Lemma 1, $\exists T\in\mathrm{GL}\left(m,\ \mathbb{F}\right)$ s.t.
+$TM=\mathrm{rref}\left(M\right)$, so $T^{-1}\mathrm{rref}\left(M\right)=M$.
+Since all except the top $r$ rows of $\mathrm{rref}\left(M\right)$ are all-zeros, $(T^{-1})_{:,:r}\mathrm{rref}\left(M\right)_{:r,:}=M$, so we can set $(C_{0},F_{0})=((T^{-1})_{:,:r},\mathrm{rref}\left(M\right)_{:r,:})$.
+Now suppose there are $C\in\mathbb{F}^{m\times r},\ F\in\mathbb{F}^{r\times n}$ s.t.
+$CF=M$.
+Since $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(CF\right)\subseteq\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)\leq r=\mathrm{rank}\left(M\right)$, we have $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)=r$.
+Since row reduction preserves rowspan, $\mathrm{rowspan}\left(\mathrm{rref}\left(M\right)\right)=\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$, so by Lemma 2, $\mathrm{rref}\left(F\right)=\mathrm{rref}\left(M\right)_{:r,:}=F_{0}$.
+Using Lemma 1, $\exists X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$ s.t.
+$XF=F_{0}$.
+Solving for $F$ and substituting it in the original equation $CF=M$ yields $(CX^{-1})F_{0}=M$.
+Since $\mathrm{rank}\left(F_{0}\right)=r$, its rows are linearly independent, so there is at most one solution for $CX^{-1}$.
+We know $C_{0}$ is one possible solution, so $CX^{-1}=C_{0}$.
+Thus, $C=C_{0}X$ and $F=X^{-1}F_{0}$.
+Since invertible matrices form a group, it follows that all pairs of full-rank factorizations of $M$ are reachable from each other via the transformation $(C,F)\mapsto(CX,X^{-1}F)$. ∎
+The upshot of this result is that there are exactly $|\mathrm{GL}\left(r,\ \mathbb{F}\right)|\in O(|\mathbb{F}|^{r^{2}})$ many rank-$r$ decompositions of a rank-$r$ matrix, no matter how large $m$ and $n$ are.
+This is the core reason why our main algorithm will run in polynomial time w.r.t.
+the size of the tensor.
+In contrast, the number of rank-$(r+1)$ decompositions of a rank-$r$ matrix $M$ might be exponential w.r.t.
+the dimensions of $M$, e.g., when $M$ is all-zeros.
+Finally, we deduce a property of matrix factorizations that are not necessarily full-rank:
+Lemma 4.
+For $U\in\mathbb{F}^{m\times r}$ and $V\in\mathbb{F}^{r\times n}$, let $r^{\prime}:=\mathrm{rank}\left(UV\right)$. Then $\mathrm{rank}\left(U\right)\leq r^{\prime}$ or $\mathrm{rank}\left(V\right)<r$.
+Proof.
+By Lemma 1, $\exists$ $T\in\mathrm{GL}\left(m,\ \mathbb{F}\right)$ s.t. $TUV=\mathrm{rref}\left(UV\right)$. In particular, all but the topmost $r^{\prime}$ rows of $TUV$ are all-zeros.
+If $\mathrm{rank}\left(V\right)=r$, then the rows of $V$ are linearly independent.
+Because $(TUV)_{i,:}=(TU)_{i,:}V=\overrightarrow{0}\ \forall i\geq r^{\prime}$, this forces $(TU)_{i,:}=\overrightarrow{0}\ \forall i\geq r^{\prime}$, so all except the top $r^{\prime}$ rows of $TU$ are all-zeros.
+We then have $\mathrm{rank}\left(TU\right)\leq r^{\prime}$ and thus $\mathrm{rank}\left(U\right)\leq r^{\prime}$ due to $T$ being invertible.
+We have thus proven that $(\mathrm{rank}\left(V\right)=r)\Rightarrow(\mathrm{rank}\left(U\right)\leq r^{\prime})$, which is logically equivalent to the final statement of the lemma. ∎
+Using the well-known fact that row-rank always equals column-rank, we can combine Lemma 4 with its transpose to get a stronger condition:
+Corollary 1.
+For $U\in\mathbb{F}^{m\times r}$ and $V\in\mathbb{F}^{r\times n}$, let $r^{\prime}:=\mathrm{rank}\left(UV\right)$.
+Then $\left(\mathrm{rank}\left(U\right)\leq r^{\prime}\ \lor\ \mathrm{rank}\left(V\right)<r\right)\ \land\ \left(\mathrm{rank}\left(U\right)<r\ \lor\ \mathrm{rank}\left(V\right)\leq r^{\prime}\right)$.
+2.2 Algorithm
+We now present our main algorithm to solve $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k$, for an arbitrary $n\times n\times n$ tensor $T$, where $R\leq 4$ is fixed.
+Our main insight is to fix $A$ and then solve for $B,C$ in a clever manner.
+To ensure we do not enumerate too many $A$, we extract a basis on the slices $T_{i,:,:}$, denoted as $\{T^{\prime}_{j,:,:}\}_{j}$, which can be obtained by flattening the slices $T_{i,:,:}$ and row-reducing the matrix $\widetilde{T}:=\begin{bmatrix}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{bmatrix}_{i}$.
+Let $J:=\mathrm{rank}\left(\widetilde{T}\right)$: then $T^{\prime}$ contains $J$ slices, and by Lemma 1, $\exists M\in\mathrm{GL}\left(n,\ \mathbb{F}\right)$ s.t.
+$\sum_{0\leq i<n}M_{j,i}T_{i,:,:}=\begin{cases}T^{\prime}_{j,:,:}&j<J\\ O&\textrm{else}\end{cases}$, implying $T_{i,:,:}=\sum_{0\leq j<J}(M^{-1})_{i,j}T^{\prime}_{j,:,:}$.
+These relations allow us to convert an arbitrary rank-$R$ decomposition of $T$ into a rank-$R$ decomposition of $T^{\prime}$ and vice versa:
+$T^{\prime}=\sum_{r=0}^{R-1}A^{\prime}_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T=\sum_{r=0}^{R-1}\left[\sum_{0\leq j<J}(M^{-1})_{i,j}A^{\prime}_{r,j}\right]_{0\leq i<n}\times B_{r,:}\times C_{r,:}$
+$T=\sum_{r=0}^{R-1}A_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T^{\prime}=\sum_{r=0}^{R-1}\left[\sum_{0\leq i<n}M_{j,i}A_{r,i}\right]_{0\leq j<J}\times B_{r,:}\times C_{r,:}$
+Thus, our original problem is equivalent to solving $T^{\prime}=\sum_{r}A^{\prime}_{r,:}\times B_{r,:}\times C_{r,:}$.
+The upshot of constructing $T^{\prime}$ is that if the original problem $T=\sum_{r}A_{r}\times B_{r,:}\times C_{r,:}$ has a solution, then every $T_{i}$ is a linear combination of $\left\{B_{r,:}\times C_{r,:}\right\}_{0\leq r<R}$, so all $T_{i,:,:}$ collectively are in a linear subspace (of $\mathbb{F}^{n\times n}$) of rank at most $R$.
+This forces $T^{\prime}$ to have at most $R$ many slices; otherwise, the original problem has no solution and we can immediately terminate.
+Thus, there are at most $|\mathbb{F}|^{R^{2}}$ possible $A^{\prime}$ to consider, which is constant w.r.t.
+$n$.
+If $J=0$, i.e. $T^{\prime}$ is empty, then $T$ must be all-zeros, so it trivially has a rank-$R$ decomposition (e.g., with all elements in $A,B,C$ set to 0).
+Otherwise, we proceed by treating each $B_{r,:}\times C_{r,:}$ as an unknown matrix $M_{r}$ with the condition that $\mathrm{rank}\left(M_{r}\right)\leq 1$ 111 Reconstructing some $B_{r,:}$ and $C_{r,:}$ from a given $M_{r}$ s.t.
+$B_{r,:}\times C_{r,:}=M_{r}$ can be done via rank factorization.
+.
+Then the decomposition problem on $T^{\prime}$ becomes a linear system of equations over matrix variables $M_{r}$ (instead of numbers): [$\ T^{\prime}_{j,:,:}\\ \vdots\end{bmatrix}_{0\leq j<J}=A^{\prime\intercal}\begin{bmatrix}\vdots\\ M_{r}\\ \vdots\end{bmatrix}_{r}$  [^2] Note that we do not concatenate the matrices $T^{\prime}_{j,:,:}$ or $M_{r}$ together; we organize them into column-vectors whose elements are themselves matrices..
+As in any linear system, we can do row reduction: by Lemma 1, $\exists F\in\mathrm{GL}\left(J,\ \mathbb{F}\right)$ s.t.
+$FA^{\prime\intercal}=\mathrm{rref}\left(A^{\prime\intercal}\right)$.
+Left-multiplying the linear system by $F$ yields [$\ D_{j}\\ \vdots\end{bmatrix}_{j}=\mathrm{rref}\left(A^{\prime\intercal}\right)\begin{bmatrix}\vdots\\ M_{r}\\ \vdots\end{bmatrix}_{r}$, where $D_{j}:=\sum_{k}F_{j,k}T^{\prime}_{k,:,:}\ \in\mathbb{F}^{n\times n}$.
+Because $F$ is invertible, we can left-multiply the new linear system by $F^{-1}$ to get the old system again, so any solution of the new system is also a solution of the old system and vice versa; thus, solving the new system over $M_{r}$ is equivalent to solving the old system.
+Now we split casework over the matrix $\mathrm{rref}\left(A^{\prime\intercal}\right)$, which we abbreviate as $E$ from this point on.
+We pay special attention to where the nonzero elements of $E$ are located within each row, specifically by analyzing the sets $S_{i}:=\left\{r:E_{i,r}\neq 0\right\}$, i.e.
+the sets of column indices of nonzero elements within each row.
+Using these sets, our system of equations is equivalent to $D_{i}=\sum_{r\in S_{i}}E_{i,r}M_{r}\ \forall i$, under the constraints $\mathrm{rank}\left(M_{r}\right)\leq 1\ \forall r$.
+2.2.1 Preprocessing
+First, we can remove any rows in $E$ that are all-zeros, since if there is such a row $j$, then having $D_{j}\neq O$ will make the system have no solution, and having $D_{j}=O$ will make this row redundant.
+Next, order for a solution to exist, we must have $\mathrm{rank}\left(D_{i}\right)\leq|S_{i}|\ \forall i$, since each $D_{i}$ is the sum of $|S_{i}|$ many matrices that each have rank $\leq 1$.
+We call this the rank precondition.
+Finally, we notice that if there is some set of indices $I$ that does not share any $M_{r}$ variables with any outside rows, i.e.
+$\forall i\in I\ \forall j\not\in I:S_{i}\cap S_{j}=\emptyset$, then solving rows $i\in I$ of the system over variables $\left\{M_{r}\right\}_{r\in\cup_{i\in I}S_{i}}$ is independent of solving rows $i\not\in I$ of the system over variables $\left\{M_{r}\right\}_{r\in\cup_{i\not\in I}S_{i}}$.
+We say $E$ is separable if this scenario is possible 333 An example of $E$ being separable is $E=\begin{bmatrix}1&&&x\\ &1\\ &&1&y\end{bmatrix}$, since row 1 can be separated from rows $\left\{0,2\right\}$.
+.
+Thus, we can assume WLOG that $E$ is a rref matrix that contains $R$ columns and is not separable 444 Algorithmically, we can separate $E$ into non-separable pieces by finding the connected components of the graph $G=\left(\left\{0,\dots J-1\right\},\ \left\{(i,j):S_{i}\cap S_{j}\neq\emptyset\right\}\right)$, which can be done in $O(R)$ time using breadth-first search.
+.
+This forces $E$ to be restricted to the following forms 555 One could further restrict $E$ by permuting the variables $M_{r}$ and rescaling them by nonzero numbers, which corresponds to right-multiplying $E$ by $PD$, where $P$ is a permutation matrix and $D$ is a diagonal matrix s.t.
+all elements on the diagonal are nonzero.
+, where $w,x,y,z$ below are arbitrary elements of $\mathbb{F}$.
+Blank elements denote zeros.
+  • $R=1$: $\begin{bmatrix}1\end{bmatrix}$
+  • $R=2$: $\begin{bmatrix}1&x\end{bmatrix}$
+  • $R=3$: $\begin{bmatrix}1&x&y\end{bmatrix},\ \begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$
+  • $R=4$: $\begin{bmatrix}1&x&y&z\end{bmatrix},\ \begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix},\ \begin{bmatrix}1&x&&y\\ &&1&z\end{bmatrix},\ \begin{bmatrix}&1&&x\\ &&1&y\end{bmatrix},\ \begin{bmatrix}1&&&x\\ &1&&y\\ &&1&z\end{bmatrix}$
+We group these forms into three types:
+2.2.2 Single row
+In this case, we only need to solve for $D_{0}$, so there exists a solution iff $\mathrm{rank}\left(D_{0}\right)\leq|S_{0}|$ (i.e.
+the rank precondition is satisfied), and such a solution can be constructed via rank factorization.
+2.2.3 Rows sharing one common column
+In this case, each pair of $S_{i}$ intersects at exactly one index, and that index is the same across all pairs, i.e.
+there is some $r^{*}$ s.t.
+for all pairs of row indices $i,i^{\prime}$ with $i\neq i^{\prime}$, $S_{i}\cap S_{i^{\prime}}=\left\{r^{*}\right\}$.
+This case covers $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix},\ \begin{bmatrix}1&x&&y\\ &&1&z\end{bmatrix},\ \begin{bmatrix}&1&&x\\ &&1&y\end{bmatrix},\ \begin{bmatrix}1&&&x\\ &1&&y\\ &&1&z\end{bmatrix}$, where all elements in the rightmost column are nonzero.
+Here, we utilize full-rank matrix factorizations.
+If $\exists i\textrm{ s.t.
+}\mathrm{rank}\left(D_{i}\right)=|S_{i}|$, we can enumerate all rank-$|S_{i}|$ factorizations of $D_{i}$, then for each factorization, check if $\mathrm{rank}\left(D_{i^{\prime}}-E_{i^{\prime},r^{*}}M_{r^{*}}\right)\leq|S_{i^{\prime}}-1|\ \forall i^{\prime}\neq i$; if this is satisfied, a solution exists, which can again be constructed with repeated matrix rank factorization.
+If there is no such $i$, then $\mathrm{rank}\left(D_{i}\right)\leq|S_{i}|-1\ \forall i$, so we can force $M_{r^{*}}=O$ and solve each $D_{i}$ independently.
+2.2.4 Rows sharing two common columns
+This case covers the last remaining form, $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$, and is the most difficult.
+Written in full, our system of equations is $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+wM_{2}+xM_{3}\\ D_{1}=&&M_{1}&+yM_{2}+zM_{3}\end{array}\right\}$, under the constraints $\mathrm{rank}\left(M_{r}\right)\leq 1\ \forall r$.
+Fixing coefficients WLOG
+WLOG assume $w,x,y,z\neq 0$, otherwise we revert to one of the previous cases (after separating $E$ into independent parts).
+We can substitute $M_{2}^{\prime}:=wM_{2},\ M_{3}^{\prime}:=xM_{3},\ y^{\prime}:=\frac{y}{w},\ z^{\prime}:=\frac{z}{x}$ to get $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}^{\prime}+M_{3}^{\prime}\\ D_{1}=&&M_{1}&+y^{\prime}M_{2}^{\prime}+z^{\prime}M_{3}^{\prime}\end{array}\right\}$, then substitute $D_{1}^{\prime}:=\frac{1}{y^{\prime}}D_{1},\ M_{1}^{\prime}:=\frac{1}{y}M_{1},\ z^{\prime\prime}:=\frac{z^{\prime}}{y^{\prime}}$ to get $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}^{\prime}+M_{3}^{\prime}\\ D_{1}^{\prime}=&&M_{1}^{\prime}&+M_{2}^{\prime}+z^{\prime\prime}M_{3}^{\prime}\end{array}\right\}$.
+Since $\mathrm{rank}\left(M\right)=\mathrm{rank}\left(\sigma M\right)$ for any $\sigma\neq 0$, solving our new system of equations with the constraint that each of $M_{0},M_{1}^{\prime},M_{2}^{\prime},M_{3}^{\prime}$ has rank $\leq 1$ is equivalent to solving our original system with the constraint that each $M_{r}$ has rank $\leq 1$.
+Thus, WLOG we can assume $w,x,y=1$ in our original system.
+If $\boldsymbol{z=1}$:
+By subtracting the second equation from the first, our system becomes $\left\{\begin{array}[]{cccc}D_{*}=&M_{0}&-M_{1}&\\ D_{1}=&&M_{1}&+M_{2}+M_{3}\end{array}\right\}$, where $D_{*}:=D_{0}-D_{1}$, so it can be solved with the rows-sharing-one-common-column case.
+Otherwise, $\boldsymbol{z\neq 0,1}$:
+We can solve several easy cases on $\mathrm{rank}\left(D_{0}\right)$ and $\mathrm{rank}\left(D_{1}\right)$.
+By the rank precondition, we must have $\mathrm{rank}\left(D_{0}\right)\leq 3$ and $\mathrm{rank}\left(D_{1}\right)\leq 3$.
+  • If $\mathrm{rank}\left(D_{0}\right)=3$:
+    We can enumerate each possible rank-3 factorization of $D_{0}$, which fixes $(M_{0},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{1}-M_{2}-zM_{3}\right)\leq 1$; if so, we have a solution.
+  • Else if $\mathrm{rank}\left(D_{1}\right)=3$:
+    Likewise, we can enumerate each rank-3 factorization of $D_{1}$, which fixes $(M_{1},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{0}-M_{2}-zM_{3}\right)\leq 1$.
+  • Else: we have $\mathrm{rank}\left(D_{0}\right)\leq 2$ and $\mathrm{rank}\left(D_{1}\right)\leq 2$:
+    • If $\mathrm{rank}\left(D_{0}\right)\leq 1$:
+      We can construct an arbitrary rank-2 factorization $D_{1}=x_{0}y_{0}+x_{1}y_{1}$, where $x_{i}$ are column-vectors and $y_{i}$ are row-vectors.
+We then try setting $M_{0}=D_{0},\ M_{1}=x_{0}y_{0},\ M_{2}+M_{3}=O,\ M_{2}+zM_{3}=x_{1}y_{1}$.
+This yields $M_{2}=-M_{3}$, so $(z-1)M_{3}=x_{1}y_{1}$.
+Since $z\neq 1$, $M_{3}=\frac{1}{z-1}x_{1}y_{1}$, so we again have a solution:
+      $\left(M_{0},M_{1},M_{2},M_{3}\right)=\left(D_{0},x_{0}y_{0},-\frac{1}{z-1}x_{1}y_{1},\frac{1}{z-1}x_{1}y_{1}\right).$
+    • Else if $\mathrm{rank}\left(D_{1}\right)\leq 1$:
+      Likewise, we can factorize $D_{0}=x_{0}y_{0}+x_{1}y_{1}$, then attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=D_{1}$ and imposing $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=O$.
+We then have the solution $M_{3}=-\frac{1}{z-1}x_{1}y_{1},\ M_{2}=-zM_{3}$.
+This leaves $(\mathrm{rank}\left(D_{0}\right),\mathrm{rank}\left(D_{1}\right))=(2,2)$ as the only remaining case.
+To solve it, we first analyze the matrices $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$, obtained by concatenating $D_{0}$ and $D_{1}$ vertically and horizontally, respectively.
+Recall that $M_{r}:=B_{r,:}\times C_{r,:}$ in our original problem; we will define column-vectors $u_{r}:=\begin{bmatrix}\vdots\\ B_{r,i}\\ \vdots\end{bmatrix}_{i}$ and row-vectors $v_{r}:=\begin{bmatrix}\cdots&C_{r,i}&\cdots\end{bmatrix}_{i}$ for brevity.
+Using the system of equations, $\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$ and $\Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]$, so a necessary condition for a solution to exist is that $\mathrm{rank}\left(\Phi\right)\leq 4$ and $\mathrm{rank}\left(\Theta\right)\leq 4$.
+We can solve two easy cases of $\mathrm{rank}\left(\Phi\right)$:
+  • If $\mathrm{rank}\left(\Phi\right)\leq 2$:
+    We can find some arbitrary rank-2 factorization $\Phi=\left[\begin{array}[]{c|c}x_{0}&x_{1}\end{array}\right]\left[\begin{array}[]{c|c}y_{0}&z_{0}\\ \hline\cr y_{1}&z_{1}\end{array}\right]$ (padding with zeros if needed) for $x_{i}\in\mathbb{F}^{n\times 1}$ and $y_{i},\ z_{i}\in\mathbb{F}^{1\times n}$; this implies $D_{0}=x_{0}y_{0}+x_{1}y_{1},\ D_{1}=x_{0}z_{0}+x_{1}z_{1}$.
+    To obtain a solution, we will attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=x_{0}z_{0}$; then we need $M_{2},M_{3}$ to satisfy $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=x_{1}z_{1}$.
+Since $x_{1}$ appears on the right-hand side of both of these equations, we will attempt further setting $u_{2}=u_{3}=x_{1}$.
+Then we have $M_{2}+M_{3}=x_{1}(v_{2}+v_{3}),\ M_{2}+zM_{3}=x_{1}(v_{2}+zv_{3})$, so we need $v_{2}+v_{3}=y_{1},\ v_{2}+zv_{3}=z_{1}$.
+Since $z\neq 1$, we can set $v_{3}=\frac{z_{1}-y_{1}}{z-1}$ and $v_{2}=y_{1}-v_{3}$.
+Thus, we can always find a solution, which written in full (in terms of $x_{i},y_{i},z_{i}$) is:
+    $\left(M_{0},M_{1},M_{2},M_{3}\right)=\left(x_{0}y_{0},x_{0}z_{0},x_{1}\left(y_{1}-\frac{z_{1}-y_{1}}{z-1}\right),x_{1}\left(\frac{z_{1}-y_{1}}{z-1}\right)\right).$
+  • Else if $\mathrm{rank}\left(\Phi\right)=4$:
+    We can enumerate all possible rank-4 factorizations of $\Phi=U_{*}W_{*}$ and check if there exist $u_{i},v_{i}$ s.t.
+$\left(\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ \left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]\right)=(U_{*},\ W_{*})$  [^6] It suffices to solve $\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, which can be done in $O(n)$ time because each $v_{i}$ is already forced to be a specific submatrix of $W_{*}$..
+By transposing the system, we can solve the cases $\mathrm{rank}\left(\Theta\right)\leq 2$ and $\mathrm{rank}\left(\Theta\right)=4$ in a similar manner.
+Thus, the remaining case we must solve is:
+If $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$ and $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$:
+We have the following matrix factorizations:
+$\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right],\ \Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right],$
+$D_{0}=\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix},\ D_{1}=\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}.$
+For brevity, we define $U:=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ W:=\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$.
+Using Corollary 1 and the property that $\mathrm{rank}\left(AB\right)\leq\mathrm{min}\left(\mathrm{rank}\left(A\right),\mathrm{rank}\left(B\right)\right)$ for any matrices $A$ and $B$, we have:
+  • $\mathrm{rank}\left(\Phi\right)=3\Rightarrow(\mathrm{rank}\left(U\right),\mathrm{rank}\left(W\right))\in\left\{(3,4),(4,3)\right\}$
+  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
+  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
+    • Since $z\neq 0$, we can replace $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}$ with $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}$ in the above statement.
+We then split casework on $\mathrm{rank}\left(U\right)$:
+  • If $\mathrm{rank}\left(U\right)=3$:
+    By rank factorization, $\exists U^{\prime}\in\mathbb{F}^{n\times 3},\ F\in\mathbb{F}^{3\times 4}$ s.t. $U=U^{\prime}F$.
+    Then $\Phi=UW=U^{\prime}(FW)$ forms a rank-3 factorization of a rank-3 matrix.
+By Lemma 3, if we can choose some arbitrary $U_{*}\in\mathbb{F}^{n\times 3},\ W_{*}\in\mathbb{F}^{3\times 2n}$ s.t.
+$\Phi=U_{*}W_{*}$ (e.g., $W_{*}=\mathrm{rref}\left(\Phi\right)$), we are guaranteed that $\exists X\in\mathrm{GL}\left(3,\ \mathbb{F}\right)$ s.t.
+$(U^{\prime},FW)=(U_{*}X^{-1},XW_{*})$.
+    Rearranging terms, we have $U=U_{*}X^{-1}F$ and $X^{-1}FW=W_{*}$.
+Defining $F^{\prime}:=X^{-1}F$ then yields $U=U_{*}F^{\prime},\ F^{\prime}W=W_{*}$.
+Since $U_{*},\ W_{*}$ are already known (as they are determined by $\Phi$) and $F^{\prime}\in\mathbb{F}^{3\times 4}$, we can enumerate all possible $F^{\prime}$ and check if $F^{\prime}W=W_{*}$ is solvable for the row-vectors $v_{i}$, which forms a linear system.
+  • Else: $\mathrm{rank}\left(U\right)=4$
+    The columns of $U$ are linearly independent, so $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right)=3$ and $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right)=3$.
+This implies $\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$ and $\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$.
+    We also have $\mathrm{rank}\left(W\right)=3$, so $\exists\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}\neq\overrightarrow{0}$ s.t.
+$\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}W=\overrightarrow{0}$, and thus:
+    $\left\{\begin{array}[]{cccc}\alpha v_{0}&&+\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ &\beta v_{1}&+\gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}.$
+    If $\alpha\neq 0$, then $v_{0}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$.
+If $\beta\neq 0$, then $v_{1}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$.
+Either case implies that $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$ is equal to the row-span of a rank-2 matrix, so $\mathrm{rank}\left(\Theta\right)\leq 2$: contradiction.
+    Thus, we need $\alpha=0$ and $\beta=0$, implying $(\gamma,\delta)\neq(0,0)$ and $\left\{\begin{array}[]{cc}\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ \gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}$.
+Subtracting between the two equations yields $(z-1)\delta v_{3}=\overrightarrow{0}$.
+Since $z\neq 1$, $\delta v_{3}=\overrightarrow{0}$, so $\gamma v_{2}=\overrightarrow{0}$.
+    Since $\gamma$ and $\delta$ are not both zero, at least one of $v_{2}$ and $v_{3}$ must be $\overrightarrow{0}$, so at least one of $M_{2}$ and $M_{3}$ is $O$.
+Either case means that the system reverts to the rows-sharing-one-common column case.
+Summary of Casework
+Below we summarize our full casework for solving the system $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}+M_{3}\\ D_{1}=&&M_{1}&+M_{2}+zM_{3}\end{array}\right\}$, where $z\neq 0,1$.
+We omit computation details for brevity, as they have already been mentioned above.
+As a reminder, we define $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$.
+  • If $\mathrm{rank}\left(\Phi\right)>4$ or $\mathrm{rank}\left(\Theta\right)>4$ or $\mathrm{rank}\left(D_{0}\right)>3$ or $\mathrm{rank}\left(D_{1}\right)>3$:
+    • No solution
+  • Else if $\mathrm{rank}\left(D_{0}\right)=3$ or $\mathrm{rank}\left(D_{1}\right)=3$:
+    • Enumerate rank factorizations of $D_{0}$ (or $D_{1}$)
+  • Else if $\mathrm{rank}\left(D_{0}\right)\leq 1$ or $\mathrm{rank}\left(D_{1}\right)\leq 1$:
+    • Set $M_{0}=D_{0}$ (or $M_{1}=D_{1}$) and use an arbitrary rank-2 factorization of $D_{1}$ (or $D_{0}$)
+  • Else if $\mathrm{rank}\left(\Phi\right)=4$ or $\mathrm{rank}\left(\Theta\right)=4$:
+    • Enumerate rank factorizations of $\Phi$ (or $\Theta$)
+  • Else if $\mathrm{rank}\left(\Phi\right)\leq 2$ or $\mathrm{rank}\left(\Theta\right)\leq 2$:
+    • Use an arbitrary rank-2 factorization of $\Phi$ (or $\Theta$)
+  • Else: we must have $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$ and $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$:
+    • Find an arbitrary rank-3 factorization $\Phi=U_{*}W_{*}$, then solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$ for $F\in\mathbb{F}^{3\times 4}$ and $v_{i}\in\mathbb{F}^{1\times n}$
+    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}\\ D_{1}=&&M_{1}&+M_{2}\end{array}\right\}$
+    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{3}\\ D_{1}=&&M_{1}&+zM_{3}\end{array}\right\}$
 2.3 Runtime
-In our algorithm, we first factor $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$ and construct $T^{\prime}$ via Gaussian elimination.
-Since $J=\mathrm{rank}\left(\overline{T}\right)$ must be at most $R$ in order for $T$ to have a rank-$R$ decomposition, we can terminate early if we detect that this condition fails, yielding a running time of $O(Rn^{3})$444 Given a $n\times m$ matrix $A$, we can construct $C,F$ s.t.
-$A=CF$, $F=\mathrm{rref}\left(A\right)$, and $C$ is invertible, as follows: 1.
-initialize $(C,F)=(I_{n},A)$ and $r=0$: 2.
-for $j$ from 00 to $m-1$ inclusive: (a) find some $i\geq r$ s.t.
-$F_{i,j}\neq 0$ (b) if such $i$ does not exist: skip to the next iteration of the for-loop (c) else: i.
-swap rows $i$ and $r$ of $F$, and swap columns $i$ and $r$ of $C$ ii.
-for $k=r+1,\dots n-1$: A.
-let $s=F_{k,j}$ B.
-set $F_{k,:}$ to $F_{k,:}-sF_{r,:}$ C.
-set $C_{:,r}$ to $C_{:,r}+sC_{:,k}$ (here $C_{:,r}$ denotes column $r$ of $C$) iii.
-increment $r$ The rank of $A$ is exactly the number of times we reach (c).
-If we only need to return $(C,F)$ under the condition $\mathrm{rank}\left(A\right)\leq R$, for some rank limit $R$, and return a failure statement otherwise, we can terminate early if we ever detect that (c) has been reached more than $R$ times.
-Thus, this procedure performs $O(Rn(m+n))$ operations.
-In our main algorithm, $A=\overline{T}$ has shape $n\times n^{2}$, so this procedure runs in $O(Rn^{3})$ time.
-.
-Then we enumerate every possible $U$, of which there are $\leq|\mathbb{F}|^{JR}$ of them.
-For each $U$, we perform Gaussian elimination on the matrix $\Xi$, which has dimensions $J\times(R+1)$ and has each element in its rightmost column being a $n\times n$ matrix, so we incur $O\left(R^{3}n^{2}\right)$ runtime.
-We then ensure there is no $r$ s.t.
-$\Omega_{r,:}=\overrightarrow{0}$ and $D_{r}\neq O$, which takes $O(J(R+n^{2}))$ time.
-If this check passes, we proceed to the rref case analysis, during which we compute the ranks of one or more matrices $D_{i}$, each of which takes $O(n^{3})$ time via Gaussian elimination.
-If $R=3$ we also have to enumerate all rank-2 decompositions of a rank-2 $n\times n$ matrix $D_{i}$ in the worst case, which takes $O(n^{3}+|\mathbb{F}|^{4}n)$ time (if we precompute all invertible $2\times 2$ matrices over $\mathbb{F}$).
-If we pass the case analysis, we construct $B$ and $C$ in $O(n^{3})$ time, then $A:=\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{r,i}$ in $O(RJn)$ time, and terminate.
-Since we are guaranteed that $J\leq R$ (as otherwise no solution exists), our final runtime is
-$O\left(Rn^{3}+\underbrace{|\mathbb{F}|^{R^{2}}}_{\textrm{fix }U}\left(\underbrace{R^{3}n^{2}}_{\textrm{row-reduce }\Xi}+\underbrace{R(n^{2}+R)}_{\textrm{check rows of }\Xi}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)+n^{3}+R^{2}n\right)\\ =O\left(|\mathbb{F}|^{R^{2}}\left(R^{3}n^{2}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)\right)$.
-Thus, for any fixed $\mathbb{F}$ and fixed $R$ our algorithm runs in $O(n^{3})$ time, with hidden constant $f(|\mathbb{F}|,R)$ polynomial in $|\mathbb{F}|$ and exponential in $R$; for a loose bound we can set $f(|\mathbb{F}|,R)=|\mathbb{F}|^{R^{2}+4}R^{3}$.
-For $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ and $R=3$, the coefficient on the $n^{3}$ term is $\approx 2^{3^{2}}\cdot(3+1)=4\cdot 2^{9}=2048$.
-2.2 Beyond rank 3
-We do not know how to make our algorithm run in $\mathrm{poly}(n)$ time for $R=4$.
-Consider the case $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$, $\mathrm{rank}\left(D_{0}\right)=2,\ \mathrm{rank}\left(D_{1}\right)=2$: unlike when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$, we cannot simply force $B_{3,:},C_{3,:}=\overrightarrow{0}$, as the system of equations might become too restrictive.
-One might also notice that $D_{0}+D_{1}=B_{0,:}\times C_{0,:}+B_{1,:}\times C_{1,:}$, so $D_{0}+D_{1}$ has rank $\leq 2$; however, if it happens to have rank 1, then there still might be exponentially many solutions for $(B_{0,:},C_{0,:},B_{1,:},C_{1,:})$, so it is too expensive to simply iterate over all such solutions and check for feasibility of $(B_{2,:},C_{2,:},B_{3,:},C_{3,:})$.
+To provide a good upper bound on the asymptotic time complexity of our algorithm, we take advantage of the fact that all matrices we deal with must have low rank in order for a solution to the original tensor decomposition problem to exist.
+For this, we utilize an efficient algorithm that returns full-rank factorizations for matrices of rank below some threshold and returns nothing otherwise.
+Lemma 5.
+Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
+  • If $\mathrm{rank}\left(M\right)\leq r^{\prime}$: return some $C\in\mathbb{F}^{m\times\mathrm{rank}\left(M\right)}$ and $F\in\mathbb{F}^{\mathrm{rank}\left(M\right)\times n}$ s.t. $M=CF$
+  • Else: return nothing
+Proof.
+We run Gaussian elimination with early termination:
+  1. 1.
+    initialize $(C,F)\leftarrow(I_{m},M)$ and $r\leftarrow 0$:
+  2. 2.
+    for $j$ from 00 to $n-1$, inclusive:
+    1. (a)
+      if there is no $r\leq i<m$ s.t. $F_{i,j}\neq 0$: continue to the next iteration of the loop
+    2. (b)
+      else:
+      1. i.
+        find some $r\leq i<m$ s.t. $F_{i,j}\neq 0$
+      2. ii.
+        swap rows $i$ and $r$ of $F$
+      3. iii.
+        swap columns $i$ and $r$ of $C$
+      4. iv.
+        let $\sigma=F_{r,j}$
+      5. v.
+        set $F_{r,:}\leftarrow F_{r,:}/\sigma$
+      6. vi.
+        set $C_{:,r}\leftarrow\sigma C_{:,r}$
+      7. vii.
+        for $k$ from $r+1$ to $m-1$, inclusive:
+        1. A.
+          let $s=F_{k,j}$
+        2. B.
+          set $F_{k,:}\leftarrow F_{k,:}-sF_{r,:}$
+        3. C.
+          set $C_{:,r}\leftarrow C_{:,r}+sC_{:,k}$
+      8. viii.
+        increment $r$
+      9. ix.
+        if $r>r^{\prime}$: terminate and return nothing
+  3. 3.
+    return $(C_{:,:r},F_{:r,:})$
+If there was no early termination (ix.), this algorithm would transform $F$ into row echelon form (not necessarily reduced row echelon form), so it would reach step (b) exactly $\mathrm{rank}\left(M\right)$ many times.
+Thus, with early termination, the algorithm will reach step (b) $\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right)$ many times, correctly terminating if $\mathrm{rank}\left(M\right)>r^{\prime}$.
+Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(n)$ time, so the total runtime is $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$.
+777 As a sidenote, we can slightly modify this algorithm to solve Lemma 1 in $O(mn\mathrm{rank}\left(M\right))$ time: remove the early termination (ix.); maintain another matrix $T$ that is initialized to $I_{m}$, and perform the same operations on it as the ones we do to $F$ throughout the algorithm;
+In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\begin{bmatrix}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{bmatrix}_{i}$.
+We also check whether $\mathrm{rank}\left(\widetilde{T}\right)\leq R$ and immediately exit if this condition is not satisfied.
+Using Lemma 5, we can complete this entire step in $O(Rn^{3})$ time.
+Then, for each possible $A^{\prime}$, we solve a system of linear equations over $R$ many matrices $M_{r}$ of rank $\leq 1$.
+Since $A^{\prime}$ has $R$ columns and $\leq R$ rows, there are $O(|\mathbb{F}|^{R^{2}})$ many of them.
+To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many.
+Then we solve each part by enumerating full-rank factorizations of a matrix with rank $\leq g(R)$, for some function $g$, and compute $O(R)$ many additional rank factorizations to extract $B_{r,:}$ and $C_{r,:}$ if a solution is found.
+To figure out $g$:
+  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, so we do not enumerate full-rank factorizations. For this, we define $g(R):=0$.
+  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2.
+  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and at least one of $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]$ or $\Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$ has rank 4.
+    • We also have the subcase where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, where we enumerate every $3\times 4$ matrix $F$ and solve for row-vectors $v_{i}$ in a linear system.
+Since this system contains $O(n)$ total scalar variables, it can be calculated in $O(n^{3})$ time, so this entire subcase can be solved in $O(|\mathbb{F}|^{12}n^{3})$ time, which is asymptotically less costly than the case where at least one of $\Phi$ or $\Theta$ has rank 4.
+Thus, we have $g(1)=0,\ g(2)=0,\ g(3)=2,\ g(4)=4$.
+Since all rank factorizations are done for $O(n)\times O(n)$ matrices, computing each one takes $O(n^{3})$ time, so we can upper bound the total time complexity of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)n^{3}\right)=O\left(\left(R+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)\right)n^{3}\right)$, which is sufficient to prove Theorem 1.
+2.3.1 Tighter analysis
+By analyzing the rank factorizations of our algorithm more carefully, we can obtain a tighter bound.
+Consider enumerating full-rank factorizations for a $O(n)\times O(n)$ matrix $M$ with rank $r$.
+We first compute an initial factorization $M=CF$, where $C$ has $r$ columns and $F$ has $r$ rows, which can be done in $O(rn^{2})$ time.
+Then, for each $X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$, we compute $(CX,X^{-1}F)$, which takes $O(nr^{2})$ time (we can precompute each $X^{-1}$ ahead of time).
+However, we may also have to set each appropriate variable matrix $M_{r}$ to a summand $(CX)_{:,r^{\prime}}(X^{-1}F)_{r^{\prime},:}$ of the rank factorization, e.g., in the rows-sharing-one-column case, in order to continue with the rest of the algorithm, so constructing these $M_{r}$ will take additional $O(rn^{2})$ time.
+Thus, the total time needed to enumerate all full-rank factorizations and set the appropriate $M_{r}$ is $O(|\mathbb{F}|^{r^{2}}(rn^{2}+r^{2}n))$.
+Lastly, we must address the one subcase under $R=4$ that we mentioned earlier: where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, and we must solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, where $F\in\mathbb{F}^{3\times 4}$, $v_{i}\in\mathbb{F}^{1\times n}$, and $W_{*}\in\mathbb{F}^{3\times 2n}$ forms the right side of a rank-3 factorization of $\Phi\in\mathbb{F}^{n\times 2n}$ (i.e.
+$\Phi=U_{*}W_{*}$ for some $U_{*}\in\mathbb{F}^{n\times 3}$).
+We can actually separate this system into $n$ many disjoint linear systems $F\begin{bmatrix}v_{0,i}&0\\ 0&v_{1,i}\\ v_{2,i}&v_{2,i}\\ v_{3,i}&zv_{3,i}\end{bmatrix}=\left[\begin{array}[]{c|c}W_{:,i}&W_{:,n+i}\end{array}\right]$ for each $0\leq i<n$, where each system only consists of $8$ variables instead of $O(n)$.
+After accounting for the enumeration of $F$ and the factorization $\Phi=U_{*}W_{*}$, we can actually solve this entire subcase in $O(3n^{2}+|\mathbb{F}|^{12}8^{3}n)$ time; this is less than the other subcase under $R=4$ where $\mathrm{rank}\left(\Phi\right)=4\ \lor\ \mathrm{rank}\left(\Theta\right)=4$, which requires $O(|\mathbb{F}|^{16}(4n^{2}+4^{2}n))$ time.
+Applying these findings to our algorithm, we enumerate rank factorizations of a $O(n)\times O(n)$ matrix of rank $g(R)$; and if a solution is found, we rank-factor the $D_{i}$ matrices, each of which must have rank $\leq R$ by the rank precondition (or else no solution exists and we terminate), which takes $O(R)\cdot O(Rn^{2})=O(R^{2}n^{2})$ additional time.
+Thus, the time complexity of our algorithm is $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}\left(g(R)n^{2}+g(R)^{2}n\right)+R^{2}n^{2}\right)\right)$.
+Notice that the $n^{3}$ now has a small coefficient compared to the other terms.
+It may be possible to obtain slightly tighter bounds through more careful analysis.
 3 With wildcards
@@
 Thus, we convert a NAE-3SAT instance with $n$ variables and $m$ clauses into a rank-2 decomposition instance on a $2n\times 2n\times 2n$ tensor with $8n+m$ fixed elements and all other elements being wildcards.
-Over arbitrary finite fields, there is an issue of scale invariance, i.e., each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t.
-$\alpha\beta\gamma=1$, allowing any one tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor.
-This scaling freedom makes constructing gadgets for NP-hardness less obvious.
+Over arbitrary finite fields, there is an issue of scale invariance: each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t.
+$\alpha\beta\gamma=1$, which allows a single tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor.
+This scaling freedom makes constructing gadgets for NP-hardness more difficult.
 3.1 Rank 1 with wildcards
-Here we outline a polynomial-time algorithm for rank-1 decomposition with wildcards over a fixed finite field $\mathbb{F}$.
-First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ and set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t.
+For rank-1 decomposition with wildcards, we give a polynomial-time algorithm over a fixed finite field $\mathbb{F}$.
+First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ of tensor indices and a set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t.
 $a_{i}b_{j}c_{k}=T_{i,j,k}\ \forall(i,j,k)\in S$.
@@
 If this check fails, there is no solution.
-Otherwise, we proceed with the observation that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve pure products, i.e., expressions of the form $\left(\prod_{i}a_{i}^{p_{i}}\right)\left(\prod_{j}b_{j}^{q_{j}}\right)\left(\prod_{k}c_{k}^{r_{k}}\right)$, for integers $p_{i},q_{j},r_{k}$.
-If we allow arbitrary multiplication and division between such expressions, all resulting expressions will still be of this form, although some $p_{i},q_{j},r_{k}$ might be negative.
-This observation suggests that we can solve this problem somewhat like a linear system of equations: construct a $|S^{\prime}|\times 3n$ integer matrix $E$ and a $|S^{\prime}|\times 1$ $\mathbb{F}$-vector $W$ s.t.
-each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ for $(i,j,k)\in S^{\prime}$ is encoded as $\prod_{m}v_{m}^{E_{l,m}}=W_{l}$ for some $l$, where the variables have been re-indexed as $v_{0,\dots n-1}=a_{0,\dots n-1},\ v_{n,\dots 2n-1}=b_{0,\dots n-1},\ v_{2n,\dots 3n-1}=c_{0,\dots n-1}$.
-Since we restrict ourselves to $S^{\prime}$, each $W_{l}$ is nonzero.
-During row reductions, elements in $W$ have to be treated differently: when adding between two rows, multiply their values in $W$; when scaling a row by $s$, raise its value in $W$ to the $s$-th power.
-However, once we reduce $\left[\begin{array}[]{c|c}E&W\end{array}\right]$ to some normal form, elements in $E$ might not be ±$\pm 1$, so extracting a solution for the variables $v_{i}$ may involve computing $r$-th roots in $\mathbb{F}$.
-An explicit example of this phenomenon is the tensor
-$T=\begin{bmatrix}\begin{bmatrix}1&?\\ ?&1\end{bmatrix},\begin{bmatrix}?&2\\ 1&?\end{bmatrix}\end{bmatrix},$
-where ?
-represents wildcards (equivalently, $S=\left\{(0,0,0),(1,1,0),(0,1,1),(1,0,1)\right\}$, $T_{0,0,0}=1,\ T_{1,1,0}=1,\ T_{0,1,1}=2,\ T_{1,0,1}=1$).
-Solving this instance is equivalent to solving the equations
-$\left\{a_{0}b_{0}c_{0}=1,\ a_{1}b_{1}c_{0}=1,\ a_{0}b_{1}c_{1}=2,\ a_{1}b_{0}c_{1}=1\right\};$
-by combining these equations, we get $\frac{(a_{0}b_{1}c_{1})(a_{1}b_{0}c_{1})}{(a_{0}b_{0}c_{0})(a_{1}b_{1}c_{0})}=\frac{2\cdot 1}{1\cdot 1}\Rightarrow\left(\frac{c_{1}}{c_{0}}\right)^{2}=2$, so solving this equation requires us to compute a square root of 2.
-To resolve this issue, we use the fact that the multiplicative group of any finite field is cyclic, so if $\mathbb{F}$ is fixed we can find a primitive element (i.e., multiplicative generator) $\mu$ via brute force and define $u_{i},w_{i}$ s.t.
-$v_{i}=\mu^{u_{i}},\ W_{i}=\mu^{w_{i}}$; then our problem reduces to solving the linear system of equations $Eu=w$ for the vector $u$ over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
+Otherwise, we observe that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve multiplications only.
+This suggests we replace all scalars with exponentials under the same base, as this turns multiplications into additions.
+Indeed, it is well-known that the multiplicative group of any finite field is cyclic, so there exists some $\mu\in\mathbb{F}$ s.t.
+its powers cover all of $\mathbb{F}\setminus\left\{0\right\}$.
+We can then substitute the variables $a_{i}=\mu^{p_{i}},\ b_{j}=\mu^{q_{j}},\ c_{k}=\mu^{r_{k}}$ and tensor elements $T_{i,j,k}=\mu^{t_{i,j,k}}$ to get the equations $p_{i}+q_{j}+r_{k}\equiv t_{i,j,k}\ (\bmod|\mathbb{F}|-1)\ \forall(i,j,k)\in S^{\prime}$.
+Thus, solving our problem is equivalent to solving a linear system over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
 From [6], solving the matrix-vector equation $xA=b$ over $\mathbb{Z}/N\mathbb{Z}$ for $n\times m$ $A$ can be done in $O(nm^{\omega-1})$ many operations, where $\omega\leq 2.371552$ [8] is the exponent for matrix multiplication, with all integers throughout the algorithm having magnitude $\leq N$.
-By transposing our system to $u^{\intercal}E^{\intercal}=w^{\intercal}$ and noting that $E$ has dimensions $|S^{\prime}|\times 3n$, we see that solving our system takes $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$.
-For fixed $\mathbb{F}$, since all integers used in this algorithm have magnitude $\leq|\mathbb{F}|-1$, each operation takes $O(1)$ time, so our algorithm runs in polynomial time.
-If we account for the number of wildcards in $T$ in the time complexity, then by letting $k$ be the number of wildcards, $|S^{\prime}|\leq|S|=n^{3}-k$, so we get $O(n(n^{3}-k)^{\omega-1})$.
+Since our system consists of $|S^{\prime}|$ equations over $3n$ variables, it can be solved in $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$.
 The main bottleneck in this algorithm is finding a primitive element of $\mathbb{F}$, for which an efficient algorithm is not known [7].
@@
 $\log(p)$ and $k$  [^9] Every element in $\mathbb{F}_{p^{k}}$ is a linear combination of $1,\alpha,\dots,\alpha^{k-1}$ with coefficients in $\mathbb{Z}/p\mathbb{Z}$, for some element $\alpha$, so every element can be specified with a $k$-element vector over $\mathbb{Z}/p\mathbb{Z}$..
-Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is w.r.t.
-input size when $p,k$ are not fixed parameters.
+Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is when $p,k$ are not fixed.
 3.2 Rank-1 with wildcards for matrices
@@
 The issue is that matrix factorizations have an immense degree of freedom, as $CF=M$ implies $(CX^{-1})(XF)=M$, so constructing gadgets seems very difficult even for rank $R=2$.
-Additionally, for rank $R=1$, we sketch an $O(n^{3})$ algorithm that does not contain $|\mathbb{F}|$ in its asymptotic running time, in contrast to the analogous problem for 3D tensors.
+Additionally, for rank $R=1$, we sketch an algorithm that runs in $O(n^{3})$ time even when $\mathbb{F}$ is not fixed (assuming each operation between elements of $\mathbb{F}$ runs in $O(1)$ time), in contrast to the analogous problem for 3D tensors.
 Again we formulate the problem as follows: given a subset $S\subseteq\left\{0,\dots n-1\right\}^{2}$ and set of elements $T_{i,j}\in\mathbb{F}\ \forall(i,j)\in S$, find variables $a_{0,\dots n-1},b_{0,\dots n-1}$ s.t.
@@
   • For $R\leq 4$, how low can the constant factor of rank-$R$ decomposition be w.r.t. to the size of the finite field and w.r.t. $R$?
-  • What is the parameterized complexity of rank-2 wildcard decomposition w.r.t.
-the number of wildcard elements in the target tensor?
-On a related note, how asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
-  • What is the parameterized complexity of rank-1 wildcard decomposition w.r.t. the size or the characteristic of the underlying finite field?
-  • Which of our algorithms be generalized beyond finite fields, such as the integers?
+  • How asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
+  • What is the parameterized complexity of rank-1 and rank-2 wildcard decomposition w.r.t. the number of wildcard elements in the target tensor?
+  • Can we have efficient tensor decomposition over number systems beyond finite fields, such as the integers?
 Acknowledgments

### v4

@@
 $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k.$
-This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the slice $[M_{r,j}]_{j}$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
+This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the $r$-th row of $M$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
 Tensor decomposition is at the heart of fast matrix multiplication, a problem that is the primary bottleneck of numerous linear algebra and graph combinatorics algorithms, such as matrix inversion and triangle detection.
@@
 Theorem 1.
-For $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for some function $f$.
-We show tighter runtime analysis in Section 2.3.1.
+For fixed $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(n^{3}+f(|\mathbb{F}|,R)n^{2})$ time, for some function $f$.
 However, when allowing certain cells of the target tensor to have arbitrary value (“wildcards”), we have a hardness result over the integers mod 2 via a reduction from Not-All-Equal 3SAT (NAE-3SAT):
@@
 We have thus proven that $(\mathrm{rank}\left(V\right)=r)\Rightarrow(\mathrm{rank}\left(U\right)\leq r^{\prime})$, which is logically equivalent to the final statement of the lemma. ∎
-Using the well-known fact that row-rank always equals column-rank, we can combine Lemma 4 with its transpose to get a stronger condition:
+Using the well-known properties $\mathrm{rank}\left(A\right)=\mathrm{rank}\left(A^{\intercal}\right)$ and $\mathrm{rank}\left(AB\right)\leq\mathrm{min}\left(\mathrm{rank}\left(A\right),\mathrm{rank}\left(B\right)\right)$ for arbitrary matrices $A$ and $B$, we can combine Lemma 4 with its transpose to get a stronger condition:
 Corollary 1.
 For $U\in\mathbb{F}^{m\times r}$ and $V\in\mathbb{F}^{r\times n}$, let $r^{\prime}:=\mathrm{rank}\left(UV\right)$.
-Then $\left(\mathrm{rank}\left(U\right)\leq r^{\prime}\ \lor\ \mathrm{rank}\left(V\right)<r\right)\ \land\ \left(\mathrm{rank}\left(U\right)<r\ \lor\ \mathrm{rank}\left(V\right)\leq r^{\prime}\right)$.
+Then $\left(\mathrm{rank}\left(U\right)\leq r^{\prime}\ \lor\ \mathrm{rank}\left(V\right)<r\right)\ \land\ \left(\mathrm{rank}\left(U\right)<r\ \lor\ \mathrm{rank}\left(V\right)\leq r^{\prime}\right)\ \land\ \mathrm{rank}\left(U\right)\geq r^{\prime}\ \land\ \mathrm{rank}\left(V\right)\geq r^{\prime}$.
 2.2 Algorithm
@@
 We can solve several easy cases on $\mathrm{rank}\left(D_{0}\right)$ and $\mathrm{rank}\left(D_{1}\right)$.
-By the rank precondition, we must have $\mathrm{rank}\left(D_{0}\right)\leq 3$ and $\mathrm{rank}\left(D_{1}\right)\leq 3$.
+WLOG assume the rank precondition: $\mathrm{rank}\left(D_{0}\right)\leq 3\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 3$.
   • If $\mathrm{rank}\left(D_{0}\right)=3$:
-    We can enumerate each possible rank-3 factorization of $D_{0}$, which fixes $(M_{0},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{1}-M_{2}-zM_{3}\right)\leq 1$; if so, we have a solution.
+    Enumerate each of the $O(|\mathbb{F}|^{9})$ possible full-rank factorizations of $D_{0}$, which fixes $(M_{0},M_{2},M_{3})$;
+    then check if $\mathrm{rank}\left(D_{1}-M_{2}-zM_{3}\right)\leq 1$.
   • Else if $\mathrm{rank}\left(D_{1}\right)=3$:
-    Likewise, we can enumerate each rank-3 factorization of $D_{1}$, which fixes $(M_{1},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{0}-M_{2}-zM_{3}\right)\leq 1$.
+    Likewise, enumerate each full-rank factorization of $D_{1}$, which fixes $(M_{1},M_{2},M_{3})$;
+    then check if $\mathrm{rank}\left(D_{0}-M_{2}-zM_{3}\right)\leq 1$.
   • Else: we have $\mathrm{rank}\left(D_{0}\right)\leq 2$ and $\mathrm{rank}\left(D_{1}\right)\leq 2$:
     • If $\mathrm{rank}\left(D_{0}\right)\leq 1$:
-      We can construct an arbitrary rank-2 factorization $D_{1}=x_{0}y_{0}+x_{1}y_{1}$, where $x_{i}$ are column-vectors and $y_{i}$ are row-vectors.
-We then try setting $M_{0}=D_{0},\ M_{1}=x_{0}y_{0},\ M_{2}+M_{3}=O,\ M_{2}+zM_{3}=x_{1}y_{1}$.
-This yields $M_{2}=-M_{3}$, so $(z-1)M_{3}=x_{1}y_{1}$.
-Since $z\neq 1$, $M_{3}=\frac{1}{z-1}x_{1}y_{1}$, so we again have a solution:
+      Set $M_{0}=D_{0}$;
+      construct an arbitrary rank-2 factorization $D_{1}=x_{0}y_{0}+x_{1}y_{1}$, where $x_{i}$ are column-vectors and $y_{i}$ are row-vectors; then set $M_{1}=x_{0}y_{0}$.
+      We now must solve $M_{2}+M_{3}=O,\ M_{2}+zM_{3}=x_{1}y_{1}$. This yields $M_{2}=-M_{3}$, and since $z\neq 1$, $M_{3}=\frac{1}{z-1}x_{1}y_{1}$, so we always have a solution:
       $\left(M_{0},M_{1},M_{2},M_{3}\right)=\left(D_{0},x_{0}y_{0},-\frac{1}{z-1}x_{1}y_{1},\frac{1}{z-1}x_{1}y_{1}\right).$
     • Else if $\mathrm{rank}\left(D_{1}\right)\leq 1$:
-      Likewise, we can factorize $D_{0}=x_{0}y_{0}+x_{1}y_{1}$, then attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=D_{1}$ and imposing $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=O$.
-We then have the solution $M_{3}=-\frac{1}{z-1}x_{1}y_{1},\ M_{2}=-zM_{3}$.
-This leaves $(\mathrm{rank}\left(D_{0}\right),\mathrm{rank}\left(D_{1}\right))=(2,2)$ as the only remaining case.
-To solve it, we first analyze the matrices $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$, obtained by concatenating $D_{0}$ and $D_{1}$ vertically and horizontally, respectively.
-Recall that $M_{r}:=B_{r,:}\times C_{r,:}$ in our original problem; we will define column-vectors $u_{r}:=\begin{bmatrix}\vdots\\ B_{r,i}\\ \vdots\end{bmatrix}_{i}$ and row-vectors $v_{r}:=\begin{bmatrix}\cdots&C_{r,i}&\cdots\end{bmatrix}_{i}$ for brevity.
-Using the system of equations, $\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$ and $\Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]$, so a necessary condition for a solution to exist is that $\mathrm{rank}\left(\Phi\right)\leq 4$ and $\mathrm{rank}\left(\Theta\right)\leq 4$.
-We can solve two easy cases of $\mathrm{rank}\left(\Phi\right)$:
-  • If $\mathrm{rank}\left(\Phi\right)\leq 2$:
-    We can find some arbitrary rank-2 factorization $\Phi=\left[\begin{array}[]{c|c}x_{0}&x_{1}\end{array}\right]\left[\begin{array}[]{c|c}y_{0}&z_{0}\\ \hline\cr y_{1}&z_{1}\end{array}\right]$ (padding with zeros if needed) for $x_{i}\in\mathbb{F}^{n\times 1}$ and $y_{i},\ z_{i}\in\mathbb{F}^{1\times n}$; this implies $D_{0}=x_{0}y_{0}+x_{1}y_{1},\ D_{1}=x_{0}z_{0}+x_{1}z_{1}$.
-    To obtain a solution, we will attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=x_{0}z_{0}$; then we need $M_{2},M_{3}$ to satisfy $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=x_{1}z_{1}$.
-Since $x_{1}$ appears on the right-hand side of both of these equations, we will attempt further setting $u_{2}=u_{3}=x_{1}$.
-Then we have $M_{2}+M_{3}=x_{1}(v_{2}+v_{3}),\ M_{2}+zM_{3}=x_{1}(v_{2}+zv_{3})$, so we need $v_{2}+v_{3}=y_{1},\ v_{2}+zv_{3}=z_{1}$.
-Since $z\neq 1$, we can set $v_{3}=\frac{z_{1}-y_{1}}{z-1}$ and $v_{2}=y_{1}-v_{3}$.
-Thus, we can always find a solution, which written in full (in terms of $x_{i},y_{i},z_{i}$) is:
-    $\left(M_{0},M_{1},M_{2},M_{3}\right)=\left(x_{0}y_{0},x_{0}z_{0},x_{1}\left(y_{1}-\frac{z_{1}-y_{1}}{z-1}\right),x_{1}\left(\frac{z_{1}-y_{1}}{z-1}\right)\right).$
-  • Else if $\mathrm{rank}\left(\Phi\right)=4$:
-    We can enumerate all possible rank-4 factorizations of $\Phi=U_{*}W_{*}$ and check if there exist $u_{i},v_{i}$ s.t.
-$\left(\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ \left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]\right)=(U_{*},\ W_{*})$  [^6] It suffices to solve $\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, which can be done in $O(n)$ time because each $v_{i}$ is already forced to be a specific submatrix of $W_{*}$..
-By transposing the system, we can solve the cases $\mathrm{rank}\left(\Theta\right)\leq 2$ and $\mathrm{rank}\left(\Theta\right)=4$ in a similar manner.
-Thus, the remaining case we must solve is:
-If $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$ and $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$:
-We have the following matrix factorizations:
-$\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right],\ \Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right],$
-$D_{0}=\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix},\ D_{1}=\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}.$
-For brevity, we define $U:=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ W:=\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$.
-Using Corollary 1 and the property that $\mathrm{rank}\left(AB\right)\leq\mathrm{min}\left(\mathrm{rank}\left(A\right),\mathrm{rank}\left(B\right)\right)$ for any matrices $A$ and $B$, we have:
-  • $\mathrm{rank}\left(\Phi\right)=3\Rightarrow(\mathrm{rank}\left(U\right),\mathrm{rank}\left(W\right))\in\left\{(3,4),(4,3)\right\}$
-  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
-  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
-    • Since $z\neq 0$, we can replace $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}$ with $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}$ in the above statement.
-We then split casework on $\mathrm{rank}\left(U\right)$:
-  • If $\mathrm{rank}\left(U\right)=3$:
-    By rank factorization, $\exists U^{\prime}\in\mathbb{F}^{n\times 3},\ F\in\mathbb{F}^{3\times 4}$ s.t. $U=U^{\prime}F$.
-    Then $\Phi=UW=U^{\prime}(FW)$ forms a rank-3 factorization of a rank-3 matrix.
-By Lemma 3, if we can choose some arbitrary $U_{*}\in\mathbb{F}^{n\times 3},\ W_{*}\in\mathbb{F}^{3\times 2n}$ s.t.
-$\Phi=U_{*}W_{*}$ (e.g., $W_{*}=\mathrm{rref}\left(\Phi\right)$), we are guaranteed that $\exists X\in\mathrm{GL}\left(3,\ \mathbb{F}\right)$ s.t.
-$(U^{\prime},FW)=(U_{*}X^{-1},XW_{*})$.
-    Rearranging terms, we have $U=U_{*}X^{-1}F$ and $X^{-1}FW=W_{*}$.
-Defining $F^{\prime}:=X^{-1}F$ then yields $U=U_{*}F^{\prime},\ F^{\prime}W=W_{*}$.
-Since $U_{*},\ W_{*}$ are already known (as they are determined by $\Phi$) and $F^{\prime}\in\mathbb{F}^{3\times 4}$, we can enumerate all possible $F^{\prime}$ and check if $F^{\prime}W=W_{*}$ is solvable for the row-vectors $v_{i}$, which forms a linear system.
-  • Else: $\mathrm{rank}\left(U\right)=4$
-    The columns of $U$ are linearly independent, so $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right)=3$ and $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right)=3$.
-This implies $\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$ and $\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$.
-    We also have $\mathrm{rank}\left(W\right)=3$, so $\exists\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}\neq\overrightarrow{0}$ s.t.
-$\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}W=\overrightarrow{0}$, and thus:
-    $\left\{\begin{array}[]{cccc}\alpha v_{0}&&+\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ &\beta v_{1}&+\gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}.$
-    If $\alpha\neq 0$, then $v_{0}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$.
-If $\beta\neq 0$, then $v_{1}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$.
-Either case implies that $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$ is equal to the row-span of a rank-2 matrix, so $\mathrm{rank}\left(\Theta\right)\leq 2$: contradiction.
-    Thus, we need $\alpha=0$ and $\beta=0$, implying $(\gamma,\delta)\neq(0,0)$ and $\left\{\begin{array}[]{cc}\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ \gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}$.
-Subtracting between the two equations yields $(z-1)\delta v_{3}=\overrightarrow{0}$.
-Since $z\neq 1$, $\delta v_{3}=\overrightarrow{0}$, so $\gamma v_{2}=\overrightarrow{0}$.
-    Since $\gamma$ and $\delta$ are not both zero, at least one of $v_{2}$ and $v_{3}$ must be $\overrightarrow{0}$, so at least one of $M_{2}$ and $M_{3}$ is $O$.
-Either case means that the system reverts to the rows-sharing-one-common column case.
-Summary of Casework
-Below we summarize our full casework for solving the system $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}+M_{3}\\ D_{1}=&&M_{1}&+M_{2}+zM_{3}\end{array}\right\}$, where $z\neq 0,1$.
-We omit computation details for brevity, as they have already been mentioned above.
-As a reminder, we define $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$.
-  • If $\mathrm{rank}\left(\Phi\right)>4$ or $\mathrm{rank}\left(\Theta\right)>4$ or $\mathrm{rank}\left(D_{0}\right)>3$ or $\mathrm{rank}\left(D_{1}\right)>3$:
-    • No solution
-  • Else if $\mathrm{rank}\left(D_{0}\right)=3$ or $\mathrm{rank}\left(D_{1}\right)=3$:
-    • Enumerate rank factorizations of $D_{0}$ (or $D_{1}$)
-  • Else if $\mathrm{rank}\left(D_{0}\right)\leq 1$ or $\mathrm{rank}\left(D_{1}\right)\leq 1$:
-    • Set $M_{0}=D_{0}$ (or $M_{1}=D_{1}$) and use an arbitrary rank-2 factorization of $D_{1}$ (or $D_{0}$)
-  • Else if $\mathrm{rank}\left(\Phi\right)=4$ or $\mathrm{rank}\left(\Theta\right)=4$:
-    • Enumerate rank factorizations of $\Phi$ (or $\Theta$)
-  • Else if $\mathrm{rank}\left(\Phi\right)\leq 2$ or $\mathrm{rank}\left(\Theta\right)\leq 2$:
-    • Use an arbitrary rank-2 factorization of $\Phi$ (or $\Theta$)
-  • Else: we must have $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$ and $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$:
-    • Find an arbitrary rank-3 factorization $\Phi=U_{*}W_{*}$, then solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$ for $F\in\mathbb{F}^{3\times 4}$ and $v_{i}\in\mathbb{F}^{1\times n}$
-    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}\\ D_{1}=&&M_{1}&+M_{2}\end{array}\right\}$
-    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{3}\\ D_{1}=&&M_{1}&+zM_{3}\end{array}\right\}$
+      Likewise, set $M_{1}=D_{1}$, factorize $D_{0}=x_{0}y_{0}+x_{1}y_{1}$, set $M_{0}=x_{0}y_{0}$.
+Then we must solve $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=O$, which can be done by setting $M_{3}=-\frac{1}{z-1}x_{1}y_{1},\ M_{2}=-zM_{3}$.
+This leaves $\mathrm{rank}\left(D_{0}\right)=2\ \land\ \mathrm{rank}\left(D_{1}\right)=2$ as the only remaining case.
+We can substitute $M_{i}=u_{i}v_{i}$, with $u_{i}\in\mathbb{F}^{n\times 1},\ v_{i}\in\mathbb{F}^{1\times n}$, since $\mathrm{rank}\left(M_{i}\right)\leq 1$.
+Then solving the system is equivalent to solving $D_{0}=\underbrace{\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]}_{=:U_{0}}\underbrace{\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]}_{=:V_{0}}$ and $D_{1}=\underbrace{\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]}_{=:U_{1}}\underbrace{\left[\begin{array}[]{c}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{array}\right]}_{=:V_{1}}$ simultaneously.
+Invoking Corollary 1, we have:
+  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow(\mathrm{rank}\left(U_{0}\right),\mathrm{rank}\left(V_{0}\right))\in\left\{(3,2),(2,3),(2,2)\right\}$
+  • $\mathrm{rank}\left(D_{1}\right)=2\Rightarrow(\mathrm{rank}\left(U_{1}\right),\mathrm{rank}\left(V_{1}\right))\in\left\{(3,2),(2,3),(2,2)\right\}$
+Collectively, we have $\left(\mathrm{rank}\left(U_{0}\right)=2\ \lor\ \mathrm{rank}\left(V_{0}\right)=2\right)\ \land\ \left(\mathrm{rank}\left(U_{1}\right)=2\ \lor\ \mathrm{rank}\left(V_{1}\right)=2\right)$.
+Although the factorizations $D_{i}=U_{i}V_{i}$ are not full-rank, at least one of $U_{i}$ and $V_{i}$ must have the same rank as $D_{i}$, which motivates the following lemma:
+Lemma 5.
+For $U\in\mathbb{F}^{m\times r},\ V\in\mathbb{F}^{r\times n}$, define $D:=UV$ and $r^{\prime}:=\mathrm{rank}\left(D\right)$.
+Let $D=U_{*}V_{*}$ be a full-rank factorization of $D$, with $U_{*}\in\mathbb{F}^{m\times r^{\prime}},\ V_{*}\in\mathbb{F}^{r^{\prime}\times n}$.
+Suppose furthermore that $\mathrm{rank}\left(U\right)=r^{\prime}$.
+Then $U=U_{*}G$ for some $G\in\mathbb{F}^{r\times r^{\prime}}$.
+Proof.
+By Lemma 3, there exists a full-rank factorization of $U=U^{\prime}F$, where $U^{\prime}\in\mathbb{F}^{m\times r^{\prime}},\ F\in\mathbb{F}^{r^{\prime}\times r}$.
+Then $D=U^{\prime}(FV)$ is a full-rank factorization of $D$, so by Lemma 3 again, $\exists X\in\mathrm{GL}\left(r^{\prime},\ \mathbb{F}\right)$ s.t.
+$(U^{\prime},FV)=(U_{*}X,X^{-1}V_{*})$, which yields $U=U_{*}(XF)$.
+666 Notice that $V$ is much less restricted, as it only has to satisfy is $(XF)V=V_{*}$, which could have exponentially many solutions w.r.t.
+$m,n$.
+Thus, we have if $UV$ is a rank-$r$ factorization of a rank-$r^{\prime}$ matrix, and that $\mathrm{rank}\left(U\right)=r$, then there are only $O(|\mathbb{F}|^{rr^{\prime}})$ possibilities for $U$, and each can be enumerated in polynomial time via matrix multiplication.
+By matrix transpose, a similar result applies to $V$ if $\mathrm{rank}\left(V\right)=r$.
+Once again, we avoid exponential dependence on the dimensions of the matrices.
+We can now solve our last case by splitting it into the following four subcases:
+  • If $\mathrm{rank}\left(U_{0}\right)=2\ \land\ \mathrm{rank}\left(U_{1}\right)=2$:
+    Enumerate all possible pairs $(U_{0},U_{1})$, which fixes the column-vectors $u_{i}$;
+    for each pair, if it does not cause conflicts on what each $u_{i}$ gets fixed to (i.e.
+if $(U_{1})_{:,2}=(U_{0})_{:,2}$ and $(U_{1})_{:,3}=(U_{0})_{:,3}$), solve the linear system $\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]$ for $v_{i}$.
+  • Else if $\mathrm{rank}\left(V_{0}\right)=2\ \land\ \mathrm{rank}\left(V_{1}\right)=2$:
+    Enumerate all possible $(V_{0},V_{1})$, which fixes the row-vectors $v_{i}$;
+    • Notice that this step relies on $z$ being nonzero, as otherwise it is possible for $v_{3}$ to still have exponentially many possibilities.
+    for each pair, if it does not cause conflicts on $v_{i}$, solve $\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$ for $u_{i}$.
+  • Else if $\mathrm{rank}\left(U_{0}\right)=2\ \land\mathrm{rank}\left(V_{1}\right)=2$:
+    Enumerate all possible $(U_{0},V_{1})$, which fixes all variables except $u_{1}$ and $v_{0}$;
+    for each pair, solve $u_{0}v_{0}=D_{0}-u_{2}v_{2}-u_{3}v_{3}$ for $v_{0}$ and $u_{1}v_{1}=D_{1}-u_{2}v_{2}-zu_{3}v_{3}$ for $u_{1}$.
+  • Else: must have $\mathrm{rank}\left(U_{1}\right)=2\ \land\mathrm{rank}\left(V_{0}\right)=2$:
+    Enumerate all possible $(U_{1},V_{0})$;
+    for each pair, solve $u_{0}v_{0}=D_{0}-u_{2}v_{2}-u_{3}v_{3}$ for $u_{0}$ and $u_{1}v_{1}=D_{1}-u_{2}v_{2}-zu_{3}v_{3}$ for $v_{1}$.
+In each subcase, when enumerating matrix pairs, we apply Lemma 5 with $r=3$ and $r^{\prime}=2$, so each matrix in the pair has $O(|\mathbb{F}|^{6})$ many possibilities, resulting in the pair having $O(|\mathbb{F}|^{12})$ many possibilities.
 2.3 Runtime
+Since $R\leq 4$, we will omit $R$ terms in some places to simplify analysis.
 To provide a good upper bound on the asymptotic time complexity of our algorithm, we take advantage of the fact that all matrices we deal with must have low rank in order for a solution to the original tensor decomposition problem to exist.
@@
 Lemma 6.
-Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
+Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(m^{2}+m(m+n)\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
   • If $\mathrm{rank}\left(M\right)\leq r^{\prime}$: return some $C\in\mathbb{F}^{m\times\mathrm{rank}\left(M\right)}$ and $F\in\mathbb{F}^{\mathrm{rank}\left(M\right)\times n}$ s.t. $M=CF$
@@
 Thus, with early termination, the algorithm will reach step (b) $\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right)$ many times, correctly terminating if $\mathrm{rank}\left(M\right)>r^{\prime}$.
-Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(n)$ time, so the total runtime is $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$.
-777 As a sidenote, we can slightly modify this algorithm to solve Lemma 1 in $O(mn\mathrm{rank}\left(M\right))$ time: remove the early termination (ix.); maintain another matrix $T$ that is initialized to $I_{m}$, and perform the same operations on it as the ones we do to $F$ throughout the algorithm;
-In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\begin{bmatrix}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{bmatrix}_{i}$.
+Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(m+n)$ time.
+Since initializing $T$ takes $O(m^{2})$ time, the total runtime is $O(m^{2}+m(m+n)\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$.
+∎
+Throughout our algorithm for tensor decomposition, we maintain the invariants that $T$ is invertible and $TM=F$. At the end of the algorithm, $F=\mathrm{rref}\left(M\right)$.
+In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\left[\begin{array}[]{c}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{array}\right]_{i}$.
 We also check whether $\mathrm{rank}\left(\widetilde{T}\right)\leq R$ and immediately exit if this condition is not satisfied.
-Using Lemma 5, we can complete this entire step in $O(Rn^{3})$ time.
+Using Lemma 6, we can complete this entire step in $O(Rn^{3})$ time, instead of the more naïve $O(n^{4})$ time.
 Then, for each possible $A^{\prime}$, we solve a system of linear equations over $R$ many matrices $M_{r}$ of rank $\leq 1$.
 Since $A^{\prime}$ has $R$ columns and $\leq R$ rows, there are $O(|\mathbb{F}|^{R^{2}})$ many of them.
-To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many.
-Then we solve each part by enumerating full-rank factorizations of a matrix with rank $\leq g(R)$, for some function $g$, and compute $O(R)$ many additional rank factorizations to extract $B_{r,:}$ and $C_{r,:}$ if a solution is found.
-To figure out $g$:
-  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, so we do not enumerate full-rank factorizations. For this, we define $g(R):=0$.
-  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2.
-  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and at least one of $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]$ or $\Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$ has rank 4.
-    • We also have the subcase where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, where we enumerate every $3\times 4$ matrix $F$ and solve for row-vectors $v_{i}$ in a linear system.
-Since this system contains $O(n)$ total scalar variables, it can be calculated in $O(n^{3})$ time, so this entire subcase can be solved in $O(|\mathbb{F}|^{12}n^{3})$ time, which is asymptotically less costly than the case where at least one of $\Phi$ or $\Theta$ has rank 4.
-Thus, we have $g(1)=0,\ g(2)=0,\ g(3)=2,\ g(4)=4$.
-Since all rank factorizations are done for $O(n)\times O(n)$ matrices, computing each one takes $O(n^{3})$ time, so we can upper bound the total time complexity of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)n^{3}\right)=O\left(\left(R+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)\right)n^{3}\right)$, which is sufficient to prove Theorem 1.
-2.3.1 Tighter analysis
-By analyzing the rank factorizations of our algorithm more carefully, we can obtain a tighter bound.
-Consider enumerating full-rank factorizations for a $O(n)\times O(n)$ matrix $M$ with rank $r$.
-We first compute an initial factorization $M=CF$, where $C$ has $r$ columns and $F$ has $r$ rows, which can be done in $O(rn^{2})$ time.
-Then, for each $X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$, we compute $(CX,X^{-1}F)$, which takes $O(nr^{2})$ time (we can precompute each $X^{-1}$ ahead of time).
-However, we may also have to set each appropriate variable matrix $M_{r}$ to a summand $(CX)_{:,r^{\prime}}(X^{-1}F)_{r^{\prime},:}$ of the rank factorization, e.g., in the rows-sharing-one-column case, in order to continue with the rest of the algorithm, so constructing these $M_{r}$ will take additional $O(rn^{2})$ time.
-Thus, the total time needed to enumerate all full-rank factorizations and set the appropriate $M_{r}$ is $O(|\mathbb{F}|^{r^{2}}(rn^{2}+r^{2}n))$.
-Lastly, we must address the one subcase under $R=4$ that we mentioned earlier: where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, and we must solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, where $F\in\mathbb{F}^{3\times 4}$, $v_{i}\in\mathbb{F}^{1\times n}$, and $W_{*}\in\mathbb{F}^{3\times 2n}$ forms the right side of a rank-3 factorization of $\Phi\in\mathbb{F}^{n\times 2n}$ (i.e.
-$\Phi=U_{*}W_{*}$ for some $U_{*}\in\mathbb{F}^{n\times 3}$).
-We can actually separate this system into $n$ many disjoint linear systems $F\begin{bmatrix}v_{0,i}&0\\ 0&v_{1,i}\\ v_{2,i}&v_{2,i}\\ v_{3,i}&zv_{3,i}\end{bmatrix}=\left[\begin{array}[]{c|c}W_{:,i}&W_{:,n+i}\end{array}\right]$ for each $0\leq i<n$, where each system only consists of $8$ variables instead of $O(n)$.
-After accounting for the enumeration of $F$ and the factorization $\Phi=U_{*}W_{*}$, we can actually solve this entire subcase in $O(3n^{2}+|\mathbb{F}|^{12}8^{3}n)$ time; this is less than the other subcase under $R=4$ where $\mathrm{rank}\left(\Phi\right)=4\ \lor\ \mathrm{rank}\left(\Theta\right)=4$, which requires $O(|\mathbb{F}|^{16}(4n^{2}+4^{2}n))$ time.
-Applying these findings to our algorithm, we enumerate rank factorizations of a $O(n)\times O(n)$ matrix of rank $g(R)$; and if a solution is found, we rank-factor the $D_{i}$ matrices, each of which must have rank $\leq R$ by the rank precondition (or else no solution exists and we terminate), which takes $O(R)\cdot O(Rn^{2})=O(R^{2}n^{2})$ additional time.
-Thus, the time complexity of our algorithm is $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}\left(g(R)n^{2}+g(R)^{2}n\right)+R^{2}n^{2}\right)\right)$.
-Notice that the $n^{3}$ now has a small coefficient compared to the other terms.
-It may be possible to obtain slightly tighter bounds through more careful analysis.
+To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many. To solve each part, we consider the most difficult subcases for each $R$:
+  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, and we must have $\mathrm{rank}\left(D_{i}\right)\leq R$ in order for a solution to exist, so we can find a solution (or prove it does not exist) in $O(n^{2})$ time.
+  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2, where we have to enumerate $O(|\mathbb{F}|^{4})$ many full-rank factorizations of such a $D_{i}$.
+For each factorization of $D_{0}$ (or $D_{1}$), we check if $D_{0}-xM_{2}$ (or $D_{1}-yM_{2}$) has rank $\leq 1$.
+Overall, solving this case takes $O(|\mathbb{F}|^{4}n^{2})$ time.
+  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and both $D_{0}$ and $D_{1}$ have rank 2, where we have to enumerate $O(|\mathbb{F}|^{12})$ many pairs of matrices in 4 subcases each.
+    Generating each pair of matrices takes $O(n)$ time, and for each pair we must solve for $X$ in some linear system $AX=B$ (or its transpose), where $A$ has shape $O(n)\times O(R)$ and $X$ has shape $O(R)\times O(n)$.
+    We can find $X$ with row reduction: find some invertible $T$ s.t.
+$TA=\mathrm{rref}\left(A\right)$ and extract a solution $X$ from the row-reduced augmented matrix $T\left[\begin{array}[]{c|c}A&B\end{array}\right]=\left[\begin{array}[]{c|c}\mathrm{rref}\left(A\right)&TB\end{array}\right]$, e.g., by checking for all-zeros rows in $\mathrm{rref}\left(A\right)$ and setting free variables of $X$ to 0.
+    We can obtain such $T$ in $O(n^{2}R)$ time by modifying the algorithm in Lemma 6: remove early termination, maintain the inverse of the $C$ matrix by performing row operations on a separate matrix, and continue the row reduction so that the $F$ matrix is transformed into rref.
+    Overall, solving this subcase takes $O(|\mathbb{F}|^{12}n^{2})$ time.
+Thus, we spend $O(|\mathbb{F}|^{g(R)}n^{2})$ time solving each system over $M_{r}$, where $g(1)=0,\ g(2)=0,\ g(3)=4,\ g(4)=12$.
+We can then upper bound the total runtime of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\cdot|\mathbb{F}|^{g(R)}n^{2}\right)$, which for fixed $R$ is equivalent to $O\left(n^{3}+|\mathbb{F}|^{R^{2}+g(R)}n^{2}\right)$.
+Notice that the large constant is multiplied with $n^{2}$, not $n^{3}$.
+It may be possible to obtain tighter bounds through more careful analysis.
 3 With wildcards